### 26 Jan 2024 v1

# Hidden Markov Models and the Bayes Filter in Categorical Probability

Abstract
We use Markov categories to develop generalizations of the theory of Markov chains and hidden Markov models in an abstract setting. This comprises characterizations of hidden Markov models in terms of local and global conditional independences as well as existing algorithms for Bayesian filtering and smoothing applicable in all Markov categories with conditionals. We show that these algorithms specialize to existing ones such as the Kalman filter, forward-backward algorithm, and the Rauch–Tung–Striebel smoother when instantiated in appropriate Markov categories. Under slightly stronger assumptions, we also prove that the sequence of outputs of the Bayes filter is itself a Markov chain with a concrete formula for its transition maps.
There are two main features of this categorical framework. The first is its generality, as it can be used in any Markov category with conditionals. In particular, it provides a systematic unified account of hidden Markov models and algorithms for filtering and smoothing in discrete probability, Gaussian probability, measure-theoretic probability, possibilistic nondeterminism and others at the same time. The second feature is the intuitive visual representation of information flow in these algorithms in terms of string diagrams.
1 Introduction
footnotetext: Acknowledgments. We thank Dario Stein for suggesting Proposition 3.2 and Matthew B. Smith for help with the forward-backward algorithm. This research was funded in part by the Austrian Science Fund (FWF) [doi:10.55776/P35992, doi:10.55776/P33122, doi:10.55776/P34129]. For open access purposes, the authors have applied a CC BY public copyright license to any author accepted manuscript version arising from this submission. AK further acknowledges funding of the Austrian Academy of Sciences (ÖAW) through the DOC scholarship 26547. YW’s research is further supported by the University of Innsbruck Early Stage Funding Program.
1.1 Context and motivation
A hidden Markov model (HMM) is a stochastic dynamical system in discrete time, whose state is considered ‘‘hidden’’ in the sense of not directly observable, together with a sequence of noisy observations that are functions of the hidden state and each point in time. HMMs are widely used across varying scientific domains to model scenarios in which the hidden state of a stochastic dynamical system must be approximately inferred from partial or imperfect observations. Making this inference can be a difficult problem in practice: while one would optimally want to employ Bayesian updating in order to make this inference, actually doing this requires conditioning on a potentially long sequence of observations, which is problematic as it requires dealing with spaces growing exponentially with the length of the sequence, so that the naïve way of calculating this quickly becomes unworkable. Recursive filters and smoothers are methods that address this difficulty by providing recursion formulas for this Bayesian update and, in some cases, employ suitable approximations that simplify things further. We limit ourselves to filters in this introduction, noting that smoothers address a more general problem by similar means, but we will treat both in the main text.
The most eminent and widely used filter is the Kalman filter, which assumes the state transitions and observations of the HMM to be linear maps with additive Gaussian noise. The most general recursive filter is the Bayes filter, which does not assume any kind of linearity, makes sense even for discrete and continuous variables alike, but is difficult to compute with in practice. But the world of recursive filters has a diverse landscape of variations upon these, each tailored to a different structure within an HMM or a different approximation scheme. Some filters are designed to handle nonlinearities in the hidden state transition and observation mappings approximately [2, 21]. Others may assume additional structure on the state and measurement spaces, such as a Riemannian metric [28]. Some may even take on a different representation of probability or information to better reflect the actual prior knowledge of the system state or improve numerical stability [4]. Some variants may also incorporate several of these features.
The differences in these variations can be a lot to keep track of. While the basic propagate-update structure is common to all of them, deriving the concrete formulas for a particular filter can be quite cumbersome. In fact, the equations that define the Kalman filter are already quite involved and difficult to gain intuition for. It would be helpful to have a unifying framework in which one can understand and reason about filters, and in which the common propagate-update structure can be made formal and precise. One may then hope that this will provide a more intuitive account of filters which moreover allows one to easily determine the relevant equations in each case.
In this paper, we aim to provide such a unifying mathematical framework in terms of categorical probability theory. Technically, we develop the theory of Markov chains, hidden Markov models, recursive Bayes filters and smoothers within any Markov category with conditionals. Indeed Markov categories serve as the stage to the actors in our paper; choosing any Markov category with conditionals makes these actors transform into statistical algorithms found in the sciences and engineering disciplines. In particular, we show that both the Kalman filter and the general Bayes filter arise upon instantiating our abstract filter construction within an appropriately chosen Markov category.
We have yet to attempt to obtain other filters as special cases of our construction, but we expect this to be possible for at least some of them, while others may require a more general framework in which approximations are considered explicitly.
1.2 Features of the categorical approach
Before outlining our abstract approach to HMMs and filters in some detail, let us list the features that distinguish it from the existing concrete constructions:
  • Generality. The abstract categorical nature of our formalism makes it applicable to a wide range of models of uncertainty. This includes discrete, continuous, and Gaussian probability, but also more exotic choices like possibilistic uncertainty [^1] As we will see in the main text, possibilistic uncertainty is not exotic at all: when applied in this context, our formalism reproduces notions familiar from automata theory. or continuous probability with an additional measurable dependence on an external parameter. In all of these cases, we therefore obtain a characterization of HMMs in terms of local or global Markov properties as well as recursion formulas for filtering and smoothing. [^2] Similar to what happens with equations for existing recursive filters, this formula may or may not be easily usable in practice. For example while the Kalman filter is easily usable and widely used, the general Bayes filter can be difficult to work with in practice with as it involves conditioning, which can be problematic especially with continuous variables.
  • Modularity. In a concrete implementation, one can implement the recursive filter as an abstract subroutine that can be instantiated automatically for each model of uncertainty. Such a model is then specified by a Markov category, which can be implemented separately, and which may also in turn be used for other algorithms besides filtering and smoothing (for example, algorithms relating to causal inference with categorical formulations [19, 10]).
    Thus, modularity facilitates code reuse both within one algorithm and across algorithms.
  • Visualization of information flow. The use of string diagrams in the categorical approach offers a convenient graphical notation, making the structure of information flow in hidden Markov models and in filtering and smoothing very explicit and intuitive to grasp. Further, the string diagrams can be easily manipulated topologically through rewrite rules. Most of the proofs in this paper involve topological manipulations of these string diagrams.
  • Probabilistic programming. Categorical probability bears close relation to the theory of probabilistic programming [42], similar to how category theory in general plays an important role in functional programming. Thanks to this connection, we expect that our approach to hidden Markov models will be particularly easy to implement in probabilistic programming languages.
1.3 Overview of the paper
We provide a summary of the main sections of the paper.
  • In Section 2, we develop the theory of hidden Markov models in Markov categories. We start this development by considering Markov chains first in Section 2.1. While traditionally, Markov chains are described measure-theoretically, or often just in discrete probability only, we extend the definition to all Markov categories with conditionals. At this level of generality, we show that the local and global Markov properties are both equivalent to the factorization property which defines a Markov chain in terms of transition kernels (Proposition 2.1). Markov chains form the basis for the rest of this paper. We then give examples of Markov chains in a few relevant Markov categories.
    Subsequently, in Section 2.2 we build hidden Markov models in a Markov category with conditionals based on the Markov chain construction. Again, we characterize these in terms of various equivalent Markov properties and in terms of a factorization into kernels (Proposition 2.8). We then give similar examples of hidden Markov models in relevant Markov categories.
  • In Section 3, we develop the categorical formulation of the Bayesian filter, which is the result of applying Bayesian inferencing to a hidden Markov model. There are several ways to compute a Bayesian filter: a given hidden Markov model can be conditioned over its observation channels in batch or recursively. We show that these constructions are equivalent (Proposition 3.2). When a fixed sequence of deterministic states is applied to these channels to act as observations, the instantiated Bayes filter results. We also show that this sequence of observations can equivalently be applied in batch or recursively (Proposition 3.7). We then give several examples of the instantiated Bayes filter in relevant Markov categories. A notable example is the Kalman filter, which is recovered from the recursive instantiated Bayes filter applied to $\mathsf{Gauss}$. We also discuss the deterministic counterparts of the Bayes filter that occur when the Markov category that forms our setting is representable. A surprising discovery to us is that the Bayesian filter itself exhibits the Markov property when recursively self-applied in a certain way. We show this in Section 3.5.
  • In Section 4, we develop the categorical formulation of the Bayesian smoother, which is very similar to the Bayesian filter but addresses the more general problem of inferring the hidden state at any time $t$ from a sequence of observations up to time $n\geq t$. We give some similar results to those obtained in Section 3, notably that there is a backward in time recursive construction of the Bayesian smoother called the fixed-interval smoother that also comes in both uninstantiated and instantiated flavors. We also develop a categorical generalization of the classic forward-backward algorithm. We then give two concrete examples of the instantiated Bayes smoother: one is to show how the forward-backward algorithm resolves to its traditional counterpart when applied in $\mathsf{FinStoch}$, and the other example demonstrates how the fixed-interval smoother recovers the Rauch-Tung-Striebel smoother when applied to $\mathsf{Gauss}$.
  • Appendix A presents the relevant background from the recent theory of Markov categories. This should make the paper accessible to anyone familiar with category theory, including symmetric monoidal categories and string diagrams, as well as basic probability. Much of the content in the appendix is summarized from [9] and [14], while the notation for conditionals is borrowed from [18].
    In more detail, we formally introduce Markov categories and illustrate the most pertinent examples. We also give several definitions, lemmas, and propositions within Markov categories that are used throughout the paper, such as almost-sure equality, determinism, representability, Kolmogorov products, and importantly, conditionals, Bayesian inversion, and conditional independence.
1.4 Detailed outline
Here, we give a brief overview of Markov categories and the main results of this paper. Traditionally, hidden Markov models are represented as graphical models corresponding to directed acyclic graphs like this:
[…]
Hence, a hidden Markov model is a particular kind of Bayesian network. The nodes of this graph are partitioned into ‘‘hidden’’ nodes (gray) and ‘‘observed’’ nodes (white) carrying the random variables $X_{i}$ and $Y_{i}$, respectively. Each directed arrow indicates a possible causal dependence between these variables. Formally, this means that a joint probability distribution $p$ is a hidden Markov model if it can be factored as
$p\big{(}x_{0},x_{1},\ldots,x_{n},y_{0},y_{1},\ldots,y_{n}\big{)}=p\mathopen{}\mathclose{{}\left(x_{0}}\right)\cdot p\mathopen{}\mathclose{{}\left(y_{0}\,|\,p_{0}}\right)\cdot\prod_{t=1}^{n}p\big{(}y_{t}\,|\,x_{t}\big{)}\cdot p\big{(}x_{t}\,|\,x_{t-1}\big{)}.$
In the categorical framework, the notion of a Bayesian network can be formulated using string diagrams, an idea that was first introduced by Fong [8] and further studied later on [35, 36, 15, 19, 10, 11]. String diagrams are a natural language for categorical probability, which is a recent high-level approach to probability theory and theoretical statistics, which abstracts away from the classical formalism of measure theory [9].
The basic constituents in categorical probability are conditional probability distributions, also known as Markov kernels. These kernels are considered as morphisms $f\colon A\to X$, drawn as boxes with $A$ as an input wire at the bottom and $X$ as an output wire on top:
[…]
Such a morphism $f$ can be thought of in two equivalent ways: first, as a process that will produce random elements of $X$ as output upon being fed any element of $A$ as input; second, as mapping every point $a\in A$ to a probability distribution on $X$ in a measurable way. In the simplest nontrivial Markov category known as $\mathsf{FinStoch}$, such an $f$ is formally given by a Markov kernel between finite sets, i.e. a stochastic matrix
$f\>:\>$ $X\times A\to\mathopen{}\mathclose{{}\left[0,1}\right]$
$\mathopen{}\mathclose{{}\left(x,a}\right)\mapsto f\mathopen{}\mathclose{{}\left(x\,|\,a}\right)$
where $f\mathopen{}\mathclose{{}\left(-\,|\,a}\right)$ is a probability distribution on $X$ for every $a\in A$. In the Markov category $\mathsf{Gauss}$, such a morphism is instead a linear map between finite-dimensional vector spaces, possibly plus Gaussian noise. In $\mathsf{BorelStoch}$, it is an arbitrary measurable Markov kernel between standard Borel measurable spaces.
If the input object $A$ is a singleton set $I$, then a morphism $p\colon I\to X$ stands for a probability distribution, depicted as a triangle like this:
[…]
So, in this situation, we do not draw $I$ in the diagram, but only one for the observation space $X$. The box $p$ represents the probability distribution, which is now considered as a map with no input and one (random) output in $X$. Therefore, we see probability distributions as special cases of the basic primitives in a categorical framework by making the input space trivial.
As the diagrammatic notation suggests, morphisms can be composed by connecting the output wire of one morphism to the input wire of another morphism. For example, a morphism $f\colon X\to Y$, representing a Markov kernel, can be composed with a morphism $p\colon I\to X$, representing a probability distribution, like this:
[…]
The result is a distribution on $Y$, which corresponds to the pushforward of $p$ along $f$. In the concrete case of $\mathsf{FinStoch}$, this composite is given by the Chapman–Kolmogorov equation
$\mathopen{}\mathclose{{}\left(f\circ p}\right)\mathopen{}\mathclose{{}\left(y}\right)\coloneqq\sum_{x\in X}f\mathopen{}\mathclose{{}\left(y\,|\,x}\right)\cdot p\mathopen{}\mathclose{{}\left(x}\right).$
Similarly, for arbitrary measurable spaces, the Chapman–Kolmogorov equation reads as
$\mathopen{}\mathclose{{}\left(f\circ p}\right)\mathopen{}\mathclose{{}\left(A}\right)\coloneqq\int_{X}f\mathopen{}\mathclose{{}\left(A\,|\,x}\right)\,p\mathopen{}\mathclose{{}\left(\mathrm{d}x}\right).$
for all $A\in\Sigma_{Y}$, where $\Sigma_{Y}$ is the $\sigma$-algebra of measurable sets in the measurable space $Y$. If the morphisms are random functions or processes, a composite $g\circ f$ can be thought of as taking an input to $f$, producing a random output, and then feeding this to $g$.
There also is a different mode of composition, often called parallel composition. This is meant in the sense of having the morphisms in question act concurrently and independently. Explicitly, given morphisms $f\colon A\to X$ and $g\colon B\to Y$, their ‘‘parallel composition’’ is a morphism $f\otimes g\colon A\otimes B\rightarrow X\otimes Y$ [^3] The operation $\otimes$” being meant in the sense of symmetric monoidal categories. depicted in terms of string diagrams as
[…]
If the morphism is a random function or process, such a parallel composite can be thought of as producing from a pair of inputs a pair of outputs by applying the two random processes independently. For example, the parallel composite of two states or distributions is the product distribution. In $\mathsf{FinStoch}$ for instance, $f\otimes g$ is explicitly defined as
$\mathopen{}\mathclose{{}\left(f\otimes g}\right)\mathopen{}\mathclose{{}\left(x,y\,|\,a,b}\right)\coloneqq f\mathopen{}\mathclose{{}\left(x\,|\,a}\right)\,g\mathopen{}\mathclose{{}\left(y\,|\,b}\right).$
To more explicitly distinguish the two notions of composition, we sometimes refer to a composite $g\circ f$ as the sequential composition while referring to a parallel composite $u\otimes v$ as the tensor (a name derived from the formalism).
A morphism with multiple inputs or outputs, depicted as
[…]
is formally a morphism $f\colon A_{1}\otimes A_{2}\otimes\dots A_{m}\rightarrow X_{1}\otimes X_{2}\otimes\dots X_{n}$. When convenient, we can group together inputs or outputs into a single wire (this can simplify things, particularly when groups of inputs/outputs are being manipulated identically). For instance, a state $p\colon I\rightarrow X\otimes Y\otimes Z$ can also be depicted with two output wires as
[…]
We will often omit wire labels when clear in context to avoid clutter (particularly for wires that are both inputs and outputs of boxes). We also often omit instances of $I$ for brevity.
Moreover, each variable can be copied or marginalized over. This is formalized in terms of special morphisms, which are drawn as string diagrams like this:
[…]
These copy and deletion maps must satisfy certain equations, which reflect elementary properties of marginalization and the formation of copying the value of a variable in probability theory, including:
  1. (i)
    If one marginalizes the output of a box, then the box and its output become irrelevant. Hence, the box can be removed from the diagram, resulting in
[…]
  2. (ii)
    If a variable gets copied and one of the outputs gets marginalized, then the resulting map acts like an identity. In terms of string diagrams, this reads as
[…]
Summarizing all relevant rules into one definition gives rise to the notion of a Markov category [9, Definition 2.1]. We recall the technical details also in Appendix A. Many different classes of probability distributions can be particular instances of Markov categories. As already indicated, there is a Markov category capturing discrete probabilities (namely $\mathsf{FinStoch}$), one for Gaussian distributions ($\mathsf{Gauss}$), one for probability theory on standard Borel spaces ($\mathsf{BorelStoch}$), and one for probability theory on arbitrary measurable spaces ($\mathsf{Stoch}$). We refer to A.4 and [9, Section 4-6] for a more detailed treatment of these instances.
Besides Markov categories modeling probability distributions, there are also non-probabilistic Markov categories. These include, for example, Markov categories in which every morphism is a deterministic map (for example $\mathsf{FinSet}$) and possibilistic Markov categories (for example $\mathsf{FinSetMulti}$). The latter is similar to $\mathsf{FinStoch}$, but differs in the fact that the outputs no longer have numerical probabilities, but are merely designated possible or impossible on every given input. In particular, a morphism is now a matrix
$f\>:\>$ $X\times A\to\mathopen{}\mathclose{{}\left\{0,1}\right\}$
$\mathopen{}\mathclose{{}\left(x,a}\right)\mapsto f\mathopen{}\mathclose{{}\left(x\,|\,a}\right)$
where $f\mathopen{}\mathclose{{}\left(x\,|\,a}\right)=1$ if output $x$ is possible on input $a$, and $f\mathopen{}\mathclose{{}\left(x\,|\,a}\right)=0$ otherwise. We refer to A.4(e) for more details.
Let us turn to the new development of this paper. In the string-diagrammatic setting, a Markov chain is represented as follows:
[…]
Here, each box $f_{i}$ represents the transition kernel which generates an $X_{i}$-valued random variable as a function of the $X_{i-1}$-valued random variable. Moreover, $X_{0}$ is directly generated from a distribution $f_{0}$ without depending on anything prior. Note that the bending of the wires has no formal meaning in this string diagram, and we use it only in order to be able to draw the diagram horizontally. Similarly, a hidden Markov model is represented as follows:
[…]
Now, additional $Y_{i}$-valued variables that depend on the $X_{i}$-valued ones are introduced. This dependence is modeled by the morphisms $g_{i}$.
Many instances of Markov categories satisfy additional properties. One relevant property for filtering of hidden Markov models is the existence of conditionals. We say that a morphism $f\colon A\to X\otimes Y$ has the conditional $f_{|X}$ if it factorizes as
[…]
In particular, every joint distribution of two random variables with values in $X$ and $Y$, possibly depending on an additional parameter from $A$, can be sampled iteratively. If every morphism in a Markov category $\mathcal{C}$ has a conditional, then we say that $\mathcal{C}$ has conditionals. Some examples of Markov categories with conditionals are $\mathsf{FinStoch}$, $\mathsf{Gauss}$, $\mathsf{BorelStoch}$ as well as $\mathsf{FinSetMulti}$. On the other hand, $\mathsf{Stoch}$, the most general Markov category for measure-theoretic probability, does not have conditionals. We refer for a more detailed treatment of conditionals as well as the closely related Bayesian inverses to A.18 and A.28 and further on to [9, Example 11.3] and [7, Definition 3.5], respectively.
For Markov categories with conditionals, we show that Markov chains and hidden Markov models can be equivalently characterized by certain conditional independence relations among the different variables, called Markov properties. This is done in Proposition 2.1 for Markov chains and in Proposition 2.8 for hidden Markov models.
Recall that a hidden Markov model often models a situation in which one only has access to observations $y_{i}\in Y_{i}$, and one wants to guess the hidden state $x_{i}\in X_{i}$ based on having recorded a sequence of values of the observations. This is the so-called filtering problem: Given values $y_{0},\ldots,y_{t}$ for the observations up to time $t$, what is our posterior belief for the value of $x_{t}$? In other words, we want to find the conditional of the hidden variable $x_{t}$ given the observations $y_{0},\ldots,y_{t}$, which by definition must be a morphism $\mathrm{BF}_{t}\colon Y_{0}\otimes\cdots\otimes Y_{t}\to X_{t}$ such that
[…]
As introduced in Section 3, this morphism $\mathrm{BF}_{t}$ is called the Bayes filter at time $t$. Here, $p_{[t]}$ is a shorthand notation for the hidden Markov model with all hidden states before time $t$ marginalized out.
Using the definition of the Bayes filter $\mathrm{BF}_{t}$ naïvely makes it challenging to work with in practice since the input space grows exponentially in $t$. This is why recursive formulas are typically used in concrete applications. In Proposition 3.2, we show that $\mathrm{BF}_{t}$ permits such a recursive formulation in every Markov category with conditionals. However, this is still not so useful in practice, where one does not need to know the Bayes filter on all inputs, but only on one particular sequence of observations $y_{0},\ldots,y_{t}$ as input. This is why we also consider the instantiated Bayes filter $\mathrm{IBF}_{t}$, which is the morphism $I\to X_{t}$ obtained by plugging in a particular sequence of observations, and prove a recursion formula for it (Proposition 3.7). Subsequently, we show that our abstract $\mathrm{IBF}_{t}$ specializes to the Kalman filter in $\mathsf{Gauss}$ (Section 3.3.4) and the discrete and continuous Bayes filters in $\mathsf{FinStoch}$ and $\mathsf{BorelStoch}$ (Section 3.3.1 and Section 3.3.2). Moreover, instantiating our general construction in $\mathsf{FinSetMulti}$ produces the possibilistic filter, which seems new.
A Markov category is called representable if every morphism $f\colon X\to Y$ can be split into a unique deterministic and a fixed sampling part. In other words, we have that
[…]
where $PY$ is an object behaving like the space of distributions on $Y$ and $f^{\sharp}\colon X\to PY$ is a deterministic morphism, which means that it commutes with copying,
[…]
and $\mathsf{samp}\colon PX\to X$ should be thought of as the process which produces a random element of $X$ distributed according to the measure that it receives as input. Examples of representable Markov categories are $\mathsf{BorelStoch}$ and $\mathsf{FinSetMulti}$.
For a hidden Markov kernel, we may think of $\mathrm{BF}_{t}^{\sharp}$ as the Bayes filter in a form that deterministically outputs the observer’s posterior distribution at time $t$ as a function of the observations. In 3.20, we show that in every representable Markov category, the deterministic Bayes filters form a Markov chain of their own, in the sense that the probability distribution in $PX_{n}$ exclusively depends on the attained distribution in $PX_{n-1}$.
A common question that is related to the filtering problem is how to incorporate future measurements into one’s guess of a hidden state. In other words, while the Bayes filter $\mathrm{BF}_{t}$ at time $t$ takes observations $y_{0},\ldots,y_{t}$ as inputs and produces an estimate of the final hidden $x_{t}$, one may also want to utilize even further observations up to $y_{n}$ for $n>t$ in order to obtain an improved estimate of $x_{t}$. This problem commonly appears in the following scenario: a Bayes filter can be used in real time as a sequence of measurements arrives in order to make a best guess of the current state using the available data at hand. Upon reaching a certain time $t=n$, one will have obtained an estimate of the state trajectory sequence $(x_{t})_{t\leq n}$, but any given estimate $x_{t}$ will only have utilized the measurements $y_{0},\ldots,y_{t}$ that came before it. At this time, one may want to ‘‘refine’’ this trajectory estimate, and obtain an improved sequence of estimates $(\hat{x}_{t})_{t\leq n}$ where each $\hat{x}_{t}$ now incorporates all measurements up to time $n$. By nature, the $n$th state estimate will not change, that is $x_{n}=\hat{x}_{n}$, but state estimates earlier in the trajectory will typically become more accurate through this process. This is the so-called smoothing problem. In Section 4, we define a Bayes smoother at time $t$ with respect to time $n$ to be a morphism $\mathrm{BS}^{n}_{t}\colon Y_{0}\otimes\cdots\otimes Y_{n}\to X_{t}$ for $t<n$ that satisfies
[…]
where we now write $p_{[n],t}$ as shorthand for the hidden Markov model up to time $n$ with all $X_{i}$ except for $X_{t}$ marginalized out.
Similar to $\mathrm{BF}_{t}$, using the definition of the Bayes smoother $\mathrm{BS}^{n}_{t}$ becomes difficult in practice as the marginalization and conditioning procedures become exponentially complex in $n$ and $t-n$. Therefore, we develop several recursive procedures, namely that in Propositions 4.4 and 4.6 and Proposition 4.9, that reduce this complexity to linear time. Again, an instantiated version of these recursions tend to be more useful in practice, so we introduce the instantiated Bayes smoother with corresponding recursions to the above propositions. We then show how our general categorical recipes specialize to the forward-backward algorithm in $\mathsf{FinStoch}$ (Section 4.3.1) and the Rauch-Tung-Striebel smoother in $\mathsf{Gauss}$ (Section 4.3.2).
Notation 1.1.
Throughout, we write
$\mathopen{}\mathclose{{}\left[n}\right]\coloneqq\mathopen{}\mathclose{{}\left\{0,1,\dots,n}\right\},\qquad\mathopen{}\mathclose{{}\left(t,n}\right]\coloneqq\mathopen{}\mathclose{{}\left\{t+1,\dots,n}\right\}$
for all $t,n\in\mathbb{N}$.
1.5 Related work
There are several different lines of work related to ours, in the sense that they also consider related notions of Markov chains, hidden Markov models, and/or filters in a categorical framework.
  • A different treatment of filters in categorical probability was developed concurrently and independently by Virgo [45]; we refer to Remark 3.5 for a discussion of the precise relation.
  • A recent line of work by McIver et al. is concerned with so-called quantitative information flow in computer science [27, 32, 1]. This seeks to establish information-theoretic upper bounds on how much certain inputs (such as sensitive secrets) of a process or program may influence certain outputs (such as ones visible to the general public). Since the hidden state of the process changes over time, it is natural to formalize this kind of situation in terms of hidden Markov models. One then looks for ways to quantify how much information the observed (leaked) values carry about the hidden state.
    These works are related to the present paper insofar as they employ a certain amount of categorical language, although everything takes place explicitly in the category of sets and stochastic matrices (the Kleisli category of the distribution monad).
  • Panangaden’s work on labelled Markov processes provides a categorical framework for Markov chains with additional inputs (called labels) rather than outputs [31]. This work places particular emphasis on developing analogues of notions from automata theory (like bisimilarity) to labelled Markov processes.
  • Recent work of Schauer and van der Meulen in [40] gives a categorical account of the Backward Filtering Forward Guiding (BFFG) algorithm for smoothing, previously introduced in [29] and further generalized in [44]. They consider a more general setting of graphical models, of which the causal structure of a hidden Markov model is an example. While they do construct the algorithm using a particular Markov category, they do not use the Markov structure explicitly. Instead, they provide a categorical treatment in terms of optics (in the sense of [16]). They then demonstrate that compatibility of the optics construction with parallel and sequential composition allows the algorithm to be formulated as a composition of optics.
In contrast to our approach, the latter three lines of work are not concerned with a general categorical formalism, but rather use categorical language to talk about the respective structures in certain particular Markov categories (roughly $\mathsf{FinStoch}$ and $\mathsf{BorelStoch}$). Also the transition kernels and the observation kernels are assumed to be the same at all times, whereas we allow them to vary over time.
2 Hidden Markov models in Markov categories
2.1 Markov chains
The goal of this subsection is to develop some basic theory of discrete-time Markov chains in Markov categories in order to set the stage for the hidden Markov models considered in the following subsection. Our main result here is that the equivalence of the various possible definitions of Markov chain: this holds not only in discrete probability, where it is most commonly considered, but also in any Markov category with conditionals (Proposition 2.1). Instantiating this result in $\mathsf{BorelStoch}$ then recovers the standard discrete case as well as the general theory of Markov chains on standard Borel spaces. Instantiating it in $\mathsf{FinSetMulti}$ specializes it to a theory of nondeterministic finite automata.
Classically, a discrete-time Markov chain is defined as a sequence of random variables [^4] We use lowercase notation for random variables since we exclusively reserve uppercase symbols for objects in Markov categories (typically sets of outcomes or measurable spaces). [^5] We here focus on the case where the Markov chain extends up to a fixed time $n\in\mathbb{N}$ only, considering the case of unbounded time in Remark 2.4. $x_{0},x_{1},\dots,x_{n}$ taking values in a common (say discrete) state space $X$, such that the sequence satisfies the Markov property
$\mathbb{P}\mathopen{}\mathclose{{}\left(x_{t}=\chi_{t}\,|\,x_{t-1}=\chi_{t-1},\dotsc,x_{0}=\chi_{0}}\right)\,=\,\mathbb{P}\mathopen{}\mathclose{{}\left(x_{t}=\chi_{t}\,|\,x_{t-1}=\chi_{t-1}}\right)$ (2.1)
for all $1\leq t\leq n$ and $\chi_{0},\dotsc,\chi_{n}\in X$ [^6] When assuming a discrete state space, this equation only needs to be considered for those sequences of values for which the left-hand side is well-defined, i.e. for which $\mathbb{P}\mathopen{}\mathclose{{}\left(x_{t-1}=\chi_{t-1},\dotsc,x_{0}=\chi_{0}}\right)>0$. A more abstract way to view this equation is to express it as the conditional independence
$x_{t}\perp x_{\mathopen{}\mathclose{{}\left[t-2}\right]}\>\mid\>x_{t-1}\qquad\forall t=1,\dotsc,n,$
where $x_{\mathopen{}\mathclose{{}\left[t-2}\right]}$ denotes the tuple of random variables $\mathopen{}\mathclose{{}\left(x_{0},\ldots,x_{t-2}}\right)$. Moreover, such a sequence is called time-homogeneous if the conditionals on the right-hand side of (2.1) are independent of $t$.
A well-known different definition of Markov chains specifies how the joint distribution is generated: one starts with an initial distribution $f_{0}$ corresponding to the law of $x_{0}$, and a transition kernel $f\colon X\to X$, then the joint distribution should be given by
$\mathbb{P}\mathopen{}\mathclose{{}\left(x_{n}=\chi_{n},\dotsc,x_{0}=\chi_{0}}\right)\,=\,f_{0}\mathopen{}\mathclose{{}\left(\chi_{0}}\right)\prod_{i=1}^{n}f\mathopen{}\mathclose{{}\left(\chi_{i}\,|\,\chi_{i-1}}\right).$
In terms of this, it can then be shown that the variables $\mathopen{}\mathclose{{}\left(x_{t}}\right)_{t\in\mathopen{}\mathclose{{}\left[n}\right]}$ form a time-homogeneous Markov chain if and only if there exist $f_{0}$ and $f$ such that the joint distributions have this form.
The following result shows an equivalence of this type for Markov chains in Markov categories (to be defined afterwards). We do not assume homogeneity in time, and we even allow the state space to vary in time. As is standard, we also consider two versions of the Markov property, a local one as above and a global one having additional conditional independences that are implied by the local ones. We refer to Section A.5 for background on conditional independence in Markov categories.
Proposition 2.1.
Let $\mathcal{C}$ be a Markov category with conditionals. Then the following are equivalent for any $p\colon I\to\bigotimes_{t\in\mathopen{}\mathclose{{}\left[n}\right]}X_{t}$ in $\mathcal{C}$:
  1. (a)
    Local Markov property: $p$ displays the conditional independence [^7] Note that the $t=1$ case of this condition is trivial because of $X_{[1-2]}=X_{\emptyset}=I$, so it would be enough to consider $t=2,\dots,n$ only.
    $X_{t}\perp X_{\mathopen{}\mathclose{{}\left[t-2}\right]}\>\mid\>X_{t-1}\qquad\forall t=1,\dotsc,n.$
  2. (b)
    Global Markov property: $p$ displays the conditional independence
    $X_{R}\perp X_{T}\>\mid\>X_{S}$
    for all disjoint sets $R,S,T\subseteq\mathopen{}\mathclose{{}\left[n}\right]$ such that for every $r\in R$ and $t\in T$ there is $s\in S$ between $r$ and $t$, meaning $r<s<t$ or $r>s>t$.
  3. (c)
    With the convention $X_{-1}\coloneqq I$, there is a sequence of morphisms ${\mathopen{}\mathclose{{}\left(f_{t}\colon X_{t-1}\to X_{t}}\right)}_{t=0}^{n}$ such that
[…]
In (2.2), the bending of the wires has no formal meaning, and we only need them to be able to draw the diagram horizontally. Our proof uses the categorical $d$-separation criterion from [10], which shows equivalences of this kind for arbitrary finite causal structures.
In the following, we also write $p_{\mathopen{}\mathclose{{}\left[t}\right]}$ for $t=0,\ldots,n$ to denote the Markov chain up to time $t$, i.e. the marginal of $p$ on $X_{0},\dots,X_{t}$. In particular, $p_{\mathopen{}\mathclose{{}\left[n}\right]}=p$.
Proof.
For (a) $\Longleftrightarrow$ (c), note that the local Markov property (a) applied to $t=n$ amounts to the existence of an $f_{n}\colon X_{n-1}\to X_{n}$ such that
[…]
Doing this recursively for all $t=n,\ldots,1$ leads to
[…]
which is Condition (c) as claimed. The converse argument is the same.
(b) $\Longleftrightarrow$ (c): Without loss of generality, we assume that $R,S,T$ form a partition of $\mathopen{}\mathclose{{}\left[n}\right]$. Then, using the $d$-separation criterion [10, Theorem 28], we have that $X_{R}\perp X_{T}\>\mid\>X_{S}$ if and only if $S$ categorically $d$-separates $R$ and $T$. This means that when removing all wires $X_{s}$ for $s\in S$, there is no connection between the wires $X_{r}$ for $r\in R$ and $X_{t}$ for $t\in T$. But this is precisely the case if for every $r\in R$ and every $t\in T$ there is $s\in S$ between $r$ and $t$. ∎
Remark 2.2.
The equivalence $(a)\Longleftrightarrow(c)$ in Proposition 2.1 can also be shown via the $d$-separation criterion [10, Theorem 28]. Due to the simplicity of the Markov chain structure, the direct argument is much simpler in this case.
Definition 2.3.
When the equivalent conditions of Proposition 2.1 hold, we say that $p$ is a Markov chain in $\mathcal{C}$. In this case:
  1. (a)
    The object $X_{t}$ is called the state space at time $t$.
  2. (b)
    The morphism $f_{0}\colon I\to X_{0}$ is the initial distribution of the chain, and the morphisms $\mathopen{}\mathclose{{}\left(f_{t}\colon X_{t-1}\to X_{t}}\right)_{t=1}^{n}$ are the transition kernels.
  3. (c)
    The initial distribution $f_{0}$ is stationary if it is preserved by the transition kernels $f_{t}$, that is if $f_{t}\circ f_{0}=f_{0}$ for all $t=1,\dots,n$. We then also call $p$ a stationary Markov chain.
  4. (d)
    We call $p$ time-homogeneous if the transition kernels $f_{t}$ can be taken to be independent of $t$ (which requires in particular the state spaces $X_{t}$ to be independent of $t$).
Remark 2.4.
Traditionally, a Markov chain is considered in unbounded time, i.e. as an infinite sequence of random variables $\mathopen{}\mathclose{{}\left(X_{t}}\right)_{t\in\mathbb{N}}$. Their joint distribution is then a probability measure on an infinite product space. Talking about this at the general categorical level poses no significant difficulties but requires an additional assumption: the existence of countable Kolmogorov products (Section A.3).
Now, if $\mathcal{C}$ is a Markov category with conditionals which also has countable Kolmogorov products, one can show an analogous equivalence as Proposition 2.1 for states
$p\colon I\to\bigotimes_{t\in\mathbb{N}}X_{t}.$
The proof is the same as the above, supplied by additional appeals to the universal property of Kolmogorov products.
Example 2.5.
In the following, we consider what Markov chains in the concrete Markov categories of A.4 look like. We exclude $\mathsf{Stoch}$ due to its lack of conditionals, while the other Markov categories under consideration do have conditionals. Concerning Kolmogorov products, $\mathsf{BorelStoch}$ is the only one among these categories which has countable Kolmogorov products, while in the other ones one can only consider Markov chains in bounded time. [^8] This is unsurprising for $\mathsf{FinStoch}$, since talking about the joint distribution of infinitely many random variables requires measure theory. Similarly for $\mathsf{Gauss}$, where considering such a joint distribution would require the consideration of Gaussians on suitable infinite-dimensional spaces.
  1. (i)
    In $\mathsf{FinStoch}$, using (2.2) shows that a Markov chain is described by the formula
    $p\mathopen{}\mathclose{{}\left(x_{0},\ldots,x_{n}}\right)=f_{0}\mathopen{}\mathclose{{}\left(x_{0}}\right)\cdot\prod_{t=1}^{n}f_{t}\mathopen{}\mathclose{{}\left(x_{t}\,|\,x_{t-1}}\right),$ (2.5)
    where $f_{0}$ is a discrete probability distribution $f_{0}\colon X_{0}\to\mathopen{}\mathclose{{}\left[0,1}\right]$ and the $f_{t}$ for $t\geq 1$ are stochastic matrices. This agrees with the usual notion of a discrete Markov chain for discrete state spaces. The $f_{t}$’s are called the transition matrices of the Markov chain [41, Section 4.1], most commonly considered in the time-homogeneous case where they are all be assumed to be equal.
  2. (ii)
    In $\mathsf{FinSetMulti}$, a Markov chain is again described by (2.5), but now with a possibility distribution $f_{0}\colon X_{0}\to\mathopen{}\mathclose{{}\left\{0,1}\right\}$ and $\mathopen{}\mathclose{{}\left\{0,1}\right\}$-valued matrices $f_{t}\colon X_{t-1}\times X_{t}\to\mathopen{}\mathclose{{}\left\{0,1}\right\}$ such that for every $x_{t}\in X_{t}$ there is at least one $x_{t-1}\in X_{t-1}$ with $f_{t}\mathopen{}\mathclose{{}\left(x_{t}\,|\,x_{t-1}}\right)=1$. Equivalently, $f_{0}$ is a subset of $X_{0}$ consisting of those elements considered as possible initial states, while $f_{t}$ is a subset of $X_{t-1}\times X_{t}$ telling us for a given state at time $t-1$ which states at time $t$ are possible. By (2.2), the joint $p\colon I\to\bigotimes_{t=0}^{n}X_{t}$ is then a subset of $\prod_{t=0}^{n}X_{t}$ consisting of those tuples $\mathopen{}\mathclose{{}\left(x_{0},\ldots,x_{n}}\right)$ which can occur as possible trajectories of the Markov chain.
    Computer scientists would consider such a time-homogeneous Markov chain a nondeterministic finite automaton. More precisely, it is a nondeterministic finite automaton with trivial input, and such that every state has at least one outgoing transition.
    Similar comments apply to a general Markov chain in $\mathsf{FinSetMulti}$, where the transition kernels are additionally allowed to depend on time.
  3. (iii)
    In $\mathsf{BorelStoch}$, a Markov chain is like in $\mathsf{FinStoch}$, but where now the objects $X_{t}$ are allowed to be arbitrary standard Borel spaces, and (2.5) now takes the form
    $p\mathopen{}\mathclose{{}\left(S_{0}\times\dots\times S_{n}}\right)=\int_{x_{0}\in S_{0}}\dots\int_{x_{n}\in S_{n}}f_{0}\mathopen{}\mathclose{{}\left(\mathrm{d}x_{0}}\right)\cdot\prod_{t=1}^{n}f_{t}\mathopen{}\mathclose{{}\left(\mathrm{d}x_{t}\,|\,x_{t-1}}\right)$
    for all measurable sets $S_{t}\subseteq X_{t}$, $t=0,\ldots,n$.
  4. (iv)
    In $\mathsf{Gauss}$, a Markov chain is given by a Gaussian vector $x_{0}$ together with relations
    $x_{t}=F_{t}x_{t-1}+\mathcal{N}\mathopen{}\mathclose{{}\left(v_{t},Q_{t}}\right)$
    for $t=1,\ldots,n$, where $F_{t}$ is a linear map and the $v_{t}$ and $Q_{t}$ are the means and covariances of independent Gaussians of the relevant dimension (which may itself depend on $t$).
Example 2.6.
Consider a time-homogeneous Markov chain in a Markov category with conditionals, given by the data of a transition kernel $f\colon X\to X$ in $\mathcal{C}$ together with a stationary distribution $f_{0}$. Then with $f_{f_{0}}^{\dagger}$ denoting the Bayesian inverse [^9] See A.28 for the definition. of $f$ with respect to $f_{0}$, meaning that
[…]
we have $f_{f_{0}}^{\dagger}\circ f_{0}=f_{0}$, so that the transition kernel $f_{f_{0}}^{\dagger}\colon X\to X$ also defines a time-homogeneous Markov chain in $\mathcal{C}$ with stationary state $f_{0}$. This is the general categorical formulation of the classical notion of the ‘‘reverse Markov chain’’ [5, Section 2.1.3], and this new Markov chain coincides with the original one if and only if the latter is reversible.
Remark 2.7.
Given a Markov category $\mathcal{C}$ and an object $W$ of $\mathcal{C}$, there is a parametric Markov category $\mathcal{C}_{W}$ associated to it [14, Section 2.2]. $\mathcal{C}_{W}$ has the same objects as $\mathcal{C}$, but the morphisms have an extra dependence on a parameter represented by $W$: a morphism $X\to Y$ in $\mathcal{C}_{W}$ is a morphism $W\otimes X\to Y$ in $\mathcal{C}$ [^10] Equivalently, $\mathcal{C}_{W}$ is the co-Kleisli category of the so-called reader comonad $W\otimes-$ on $\mathcal{C}$. It is also known that if $\mathcal{C}$ has conditionals, then so does $\mathcal{C}_{W}$ [14, Lemma 2.10].
We can thus apply the machinery of this section to $\mathcal{C}_{W}$ as well, and this produces automatically a notion of a ‘‘Markov chain parametrized by $W$’’, which is now (for $n$ time steps) represented in the original category $\mathcal{C}$ by a morphism $p\colon W\to\bigotimes_{t=0}^{n}X_{t}$. The equivalence between the Markov properties and the description in terms of initial distribution and transition kernels is thus an instance of Proposition 2.1 as well, since $\mathcal{C}_{W}$ is a Markov category with conditionals. The initial distribution and transition kernels are represented in $\mathcal{C}$ by $f_{0}\colon W\to X_{0}$ and $f_{t}\colon W\otimes X_{t-1}\to X_{t}$, respectively. The joint distribution $p$ corresponds to the string diagram which in $\mathcal{C}$ takes the shape
[…]
This illustrates again that $W$ can be thought of as an ‘‘external’’ parameter that is fed into each of the transition kernels.
2.2 Hidden Markov models
We now turn to hidden Markov models, for which the idea is that one has a Markov chain of ‘‘hidden states’’, as in the previous section, together with a sequence of ‘‘observations’’ $g_{t}\colon X_{t}\to Y_{t}$ that depend on the hidden states. The joint distribution of these observations is typically not a Markov chain in its own right since it need not fulfill the local Markov property: given the observation at time $t$, learning about an observation at an earlier time can help us make a better guess about what the observation at time $t+1$ will be.
The following result will serve as the definition of a hidden Markov model in a Markov category. We abbreviate $\bigotimes_{t\in\mathopen{}\mathclose{{}\left[n}\right]}\mathopen{}\mathclose{{}\left(X_{t}\otimes Y_{t}}\right)$ to $\mathopen{}\mathclose{{}\left(X\otimes Y}\right)_{\mathopen{}\mathclose{{}\left[n}\right]}$.
Proposition 2.8.
Let $\mathcal{C}$ be a Markov category with conditionals, and let $\mathopen{}\mathclose{{}\left(X_{t}}\right)_{t\in\mathopen{}\mathclose{{}\left[n}\right]}$ and $\mathopen{}\mathclose{{}\left(Y_{t}}\right)_{t\in\mathopen{}\mathclose{{}\left[n}\right]}$ be two families of objects. Then the following are equivalent for any state $p\colon I\to\mathopen{}\mathclose{{}\left(X\otimes Y}\right)_{\mathopen{}\mathclose{{}\left[n}\right]}$:
  1. (i)
    The backward Markov property: for all $t=1,\dots,n$, we have
    1. (a)
      $X_{t}\perp X_{\mathopen{}\mathclose{{}\left[t-2}\right]},Y_{\mathopen{}\mathclose{{}\left[t-1}\right]}\,|\,X_{t-1}\qquad\forall t=1,\dots,n$,
    2. (b)
      $Y_{t}\perp Y_{\mathopen{}\mathclose{{}\left[t-1}\right]},X_{\mathopen{}\mathclose{{}\left[t-1}\right]}\,|\,X_{t}\hskip 36.0pt\forall t=1,\dots,n$.
  2. (ii)
    The local Markov property: for all $t=1,\dots,n$, we have
    1. (a)
      $X_{t}\perp X_{\mathopen{}\mathclose{{}\left[t-2}\right]},Y_{\mathopen{}\mathclose{{}\left[t-1}\right]}\,|\,X_{t-1}\qquad\forall t=1,\dots,n$,
    2. (b)
      $Y_{t}\perp X_{\mathopen{}\mathclose{{}\left[n}\right]\setminus\mathopen{}\mathclose{{}\left\{t}\right\}},Y_{\mathopen{}\mathclose{{}\left[n}\right]\setminus\mathopen{}\mathclose{{}\left\{t}\right\}}\,|\,X_{t}\qquad\forall t=0,\dots,n$.
  3. (iii)
    The global Markov property: for all disjoint sets $R,S,T\subseteq\mathopen{}\mathclose{{}\left[n}\right]$ and disjoint sets $U,V,W\subseteq\mathopen{}\mathclose{{}\left[n}\right]$ such that
    $\forall r\in R,\,t\in T\;\,\exists s\in S\colon s\text{ is between }r\text{ and }t$
    and
    $\forall u\in U,\,w\in W\;\,\exists s\in S\colon s\text{ is non-strictly between }u\text{ and }w$
    we have
    $X_{R},Y_{U}\perp X_{T},Y_{W}\,|\,X_{S},Y_{V}.$
  4. (iv)
    There exist a state $f_{0}\colon I\to X_{0}$ and sequences of morphisms
    $f_{t}\colon X_{t-1}\to X_{t},\qquad g_{t}\colon X_{t}\to Y_{t},\qquad t=0,\dots,n$
    such that
[…]
These Markov properties are surprisingly difficult to find in the traditional literature on hidden Markov models. The backward Markov property is stated, for example, in [26], while two special cases of the global Markov property appear at [5, Corollary 2.2.5]. But we are not aware of any reference proving the above equivalences in measure-theoretic probability.
Notation 2.9.
In the following, we write:
  1. (a)
    $p_{\mathopen{}\mathclose{{}\left[t}\right]}$ for the marginal of $p$ up to time $t$, i.e. the marginal on $\mathopen{}\mathclose{{}\left(X\otimes Y}\right)_{\mathopen{}\mathclose{{}\left[t}\right]}$.
  2. (b)
    $p^{X}$ for the marginal of $p$ on $X_{\mathopen{}\mathclose{{}\left[n}\right]}$, and similarly $p^{Y}$ for the marginal of $p$ on $Y_{\mathopen{}\mathclose{{}\left[n}\right]}$.
Both notations can also be combined to $p_{\mathopen{}\mathclose{{}\left[t}\right]}^{X}$ and $p_{\mathopen{}\mathclose{{}\left[t}\right]}^{Y}$.
Proof.
First, the global Markov property (iii) trivially implies the local Markov property (ii) by making suitable choices of the sets involved.
We next show that the local Markov property (ii) implies the backward Markov property (i). Their respective conditions (i)(a) already coincide. Condition (i)(b) of the backward Markov property follows from Condition (ii)(b) of the local Markov property by marginalizing out $X_{\mathopen{}\mathclose{{}\left[n}\right]\setminus\mathopen{}\mathclose{{}\left[t}\right]}$ and $Y_{\mathopen{}\mathclose{{}\left[n}\right]\setminus\mathopen{}\mathclose{{}\left[t}\right]}$, which preserves conditional independence [9, Lemma 12.5(b)].
We now show that the backward Markov property (i) is sufficient for $p$ to have the form (2.7). Applying Condition (i)(b) of the backward Markov property for $t=n$ gives
[…]
for suitable $g_{n}$ by Proposition A.30(b), where we leave it understood that the wires have been reordered to match the codomain to be $X_{\mathopen{}\mathclose{{}\left[n-1}\right]}\otimes Y_{\mathopen{}\mathclose{{}\left[n-1}\right]}\otimes X_{n}\otimes Y_{n}$ rather than $\mathopen{}\mathclose{{}\left(X\otimes Y}\right)_{\mathopen{}\mathclose{{}\left[n}\right]}$. Similarly, we apply Condition (i)(a) to obtain $f_{n}$ such that
[…]
Combining the previous two equations and applying the same procedure to $p_{\mathopen{}\mathclose{{}\left[n-1}\right]}$ iteratively produces the desired (2.7).
Finally, we show that every hidden Markov model in the form (2.7) satisfies the global Markov property (iii). To this end, we apply the $d$-separation criterion [10, Theorem 34] to the causal structure (2.7). As in the proof of Proposition 2.1, this causal structure is formalized as a cospan of hypergraphs, but we omit these formal details now. In order to apply the categorical $d$-separation criterion [10, Definition 19], let
$R+U\subseteq\mathopen{}\mathclose{{}\left[n}\right]+\mathopen{}\mathclose{{}\left[n}\right],\qquad S+V\subseteq\mathopen{}\mathclose{{}\left[n}\right]+\mathopen{}\mathclose{{}\left[n}\right],\qquad T+W\subseteq\mathopen{}\mathclose{{}\left[n}\right]+\mathopen{}\mathclose{{}\left[n}\right]$
be three disjoint sets of output wires, with the first component of each set of type $X$ and the second of type $Y$. Then $R+U$ is $d$-separated from $T+W$ by $S+V$ if and only if after marginalizing out all wires not contained in either set, and removing all wires in $S+V$, the sets $R+U$ and $T+W$ become disconnected. Using formal reasoning similar to the one in the proof of Proposition 2.1, it becomes clear that this happens if and only if the betweenness condition of (iii) is satisfied. Hence, the global Markov property in the general sense [10, Definition 27] specializes to the global Markov property (iii), as desired. ∎
The following terminology is now in addition to the one for Markov chains from 2.3.
Definition 2.10.
When the equivalent conditions of Proposition 2.8 hold, we say that $p$ is a hidden Markov model in $\mathcal{C}$. In this case:
  1. (a)
    The object $Y_{t}$ is called the observation space at time $t$.
  2. (b)
    The morphisms $\mathopen{}\mathclose{{}\left(g_{t}\colon X_{t}\to Y_{t}}\right)_{t=0,\dots,n}$ are the observation kernels.
  3. (c)
    We call $p$ time-homogeneous if both the transition kernels $f_{t}$ and the observation kernels $g_{t}$ can be taken to be independent of $t$ (which requires, in particular, the state spaces $X_{t}$ and observation spaces $Y_{t}$ to be independent of $t$).
So, most concretely, we can say that a hidden Markov model in $\mathcal{C}$ is specified by a Markov chain ${\mathopen{}\mathclose{{}\left(f_{t}\colon X_{t-1}\to X_{t}}\right)}_{t=1}^{n}$ with initial distribution $f_{0}\colon I\to X_{0}$ and an additional sequence of morphisms $\mathopen{}\mathclose{{}\left(g_{t}\colon X_{t}\to Y_{t}}\right)_{t=0}^{n}$.
Remark 2.11.
As in the case of Markov chains of Proposition 2.1, one can prove in Markov categories with countable Kolmogorov products a similar version for morphisms $p\colon I\to\mathopen{}\mathclose{{}\left(X\otimes Y}\right)_{\mathbb{N}}$ by using the universal property of Kolmogorov products to rewrite (2.7) equivalently as a sequence of equations for the marginals up to time $n$.
Remark 2.12.
Many authors assume in the definition of a hidden Markov model that both the state spaces $X_{t}$ and the observation spaces $Y_{t}$ are independent of $t$, and that the observations are obtained from the hidden states via a fixed Markov kernel $g\colon X\to Y$. This is a form of time-homogeneity for hidden Markov models.
Example 2.13.
In the following, we work out more explicitly what hidden Markov models are in the concrete Markov categories of A.4.
  1. (i)
    In $\mathsf{FinStoch}$, a hidden Markov model is described by the formula
    $p\mathopen{}\mathclose{{}\left(x_{0},y_{0},\ldots,x_{n},y_{n}}\right)=\prod_{t=0}^{n}f_{t}\mathopen{}\mathclose{{}\left(x_{t}\,|\,x_{t-1}}\right)\cdot g_{t}\mathopen{}\mathclose{{}\left(y_{t}\,|\,x_{t}}\right)$ (2.9)
    where $x_{-1}$ is the only element of the singleton set $X_{-1}\coloneqq I$, and the $f_{t}$ and $g_{t}$ are stochastic matrices as usual. The $g_{t}$ are also called emission probabilities.
  2. (ii)
    For $\mathsf{FinSetMulti}$, we already saw in 2.5 that a Markov chain corresponds to a nondeterministic finite automaton with a trivial input alphabet (and without termination). The additional morphisms $g_{t}\colon X_{t}\to Y_{t}$ can then be considered as output maps of the automaton, which assign a nonempty set of possible outputs to every state. Like this, a string $y_{0},\ldots,y_{n}$ is a possible sequence of observations for $p$ if and only if there exists a possible string $x_{0},\ldots,x_{n}$ with respect to the underlying Markov chain, and additionally such that $y_{t}$ is possible given $x_{t}$ according to $g_{t}$. Such automata are closely related to labelled transition systems, provided that one thinks of the labels as the outputs of the automaton. One obtains the closest match between a labelled transition system and a notion of a hidden Markov model if one replaces the transition and observation kernels with a single sequence of maps
    $h_{t}\colon X_{t-1}\longrightarrow X_{t}\times Y_{t},$
    which one may think of as outputting a pair consisting of an updated state and an observation at every time $t$. This corresponds to our case if these maps have the specific form
[…]
    but in general the observation at time $t$ may depend on the previous state as well. [^11] We could develop the theory of hidden Markov models to match this type of setting, which is sometimes done in the classical literature, e.g. in [6]. However, this notion is not significantly more general, and we have decided to generalize the more commonly considered notion of hidden Markov models to the categorical setting. So this generalized notion of hidden Markov models would instantiate to labelled transition systems in $\mathsf{FinSetMulti}$, with the additional requirement that every state must have at least one outgoing (labelled) transition. Our version of 2.10 instantiates in $\mathsf{FinStoch}$ to a more specific kind of labelled transition system.
  3. (iii)
    In $\mathsf{Gauss}$, a hidden Markov model is given by a Gaussian vector $x_{0}$ together with relations
    $x_{t}=A_{t}x_{t-1}+\mathcal{N}\mathopen{}\mathclose{{}\left(v_{t},Q_{t}}\right),\qquad y_{t}=H_{t}x_{t}+\mathcal{N}\mathopen{}\mathclose{{}\left(w_{t},R_{t}}\right),$
    where the $F_{t}$ and $G_{t}$ are matrices of the relevant sizes, and $x_{-1}=0$ because of $X_{-1}=I=\mathbb{R}^{0}$. Hence $x_{t}$ is an affine function of $x_{t-1}$ with offset given by $v_{t}$ plus a centered Gaussian noise, and similarly each observation $y_{t}$ is such a function of the state $x_{t}$. This description perfectly matches the existing notion of Gaussian hidden Markov models [38].
  4. (iv)
    A hidden Markov model in $\mathsf{BorelStoch}$ is similar to one in $\mathsf{FinStoch}$. However, the state and observation spaces can now be arbitrary standard Borel spaces, and the transition and observation kernels can be arbitrary measurable Markov kernels. It is worth emphasizing again that the equivalence with the Markov properties from Proposition 2.8 still holds in this case and just as easy to prove as in $\mathsf{FinStoch}$, while a traditional approach would easily get lost in measure-theoretic details.
Remark 2.14.
As in Remark 2.7, we can speak of hidden Markov models in the parametric Markov category $\mathcal{C}_{W}$ associated to any Markov category $\mathcal{C}$ with conditionals and an object $W$. As with the case of Markov chains, the data of such a hidden Markov model in $\mathcal{C}_{W}$ can be entirely translated to that of its representatives in $\mathcal{C}$, which amounts to the initial distribution, transition kernels and observation kernels all having a possible extra dependence on $W$. This results in decomposition as in (2.7), where now a copy of the input $W$ is fed into all of the $f_{t}$ and $g_{t}$.
3 The Bayes filter in Markov categories
Throughout this section, we work in a Markov category $\mathcal{C}$ with conditionals and consider a hidden Markov model in the notation of 2.10. As mentioned, the idea is to model situations in which a system evolves in a Markovian fashion, but the states of the system are only indirectly observable through noisy observations. A natural problem to ask now is, given a sequence of observations $\mathopen{}\mathclose{{}\left(y_{i}}\right)_{i=0}^{t}$ up to a time $t$, with $y_{i}\in Y_{i}$, what can we infer about the hidden state $x_{t}\in X_{t}$? This is precisely the problem of filtering.
3.1 The Bayes filter
Using Bayesian inference, answering the filtering question amounts to conditioning $x_{t}$ with respect to the given sequence of observations $y_{0},\dotsc,y_{t}$. Here is the formulation in our categorical setting.
Definition 3.1.
For $p\colon I\to\mathopen{}\mathclose{{}\left(Y\otimes X}\right)_{\mathopen{}\mathclose{{}\left[n}\right]}$ a hidden Markov model, the Bayes filter at time $t\in\mathopen{}\mathclose{{}\left[n}\right]$ is the conditional
[…]
Note that the marginalizations on the right-hand side are precisely those over the state spaces $X_{0},\dots,X_{t-1}$. We can therefore also phrase this definition in the following form: the Bayes filter at time $t$ is any morphism $\mathrm{BF}_{t}\colon Y_{\mathopen{}\mathclose{{}\left[t}\right]}\to X_{t}$ such that
[…]
Since the definition of $\mathrm{BF}_{t}$ is difficult to work with in practice due to the space $Y_{\mathopen{}\mathclose{{}\left[t}\right]}$ that we condition on having size exponential in $t$ (when it is finite), we now develop recursion formulas for this conditional which generalize known formulas to the categorical setting.
Proposition 3.2.
The Bayes filter can be recursively computed through conditionals as [^12] This formulation was suggested to us by Dario Stein.
[…]
for all $t=0,\dots,n$, where we take $\mathrm{BF}_{-1}\coloneqq\mathrm{id}_{I}$ to be the trivial morphism.
Setting $\mathrm{BF}_{-1}\coloneqq\mathrm{id}_{I}$ makes sense insofar as before time 00, there are no observations and no hidden state. Hence, the Bayes filter at that time takes no input and produces no output. Plugging this in explicitly amounts to the equation
[…]
so that $\mathrm{BF}_{0}$ is a Bayesian inverse $\mathopen{}\mathclose{{}\left(g_{0}}\right)_{f_{0}}^{\dagger}$ of $g_{0}$ with respect to $f_{0}$ (in the sense of A.28).
Proof.
Recall the definition
[…]
Now we expand $p_{\mathopen{}\mathclose{{}\left[t}\right]}$ via (2.7) and use Proposition A.24(a) to split this into two conditionals,
{tikzpicture}${\begin{tikzpicture}}\quad=_{{p_{\mathopen{}\mathclose{{}\left[t}\right]}^{Y}\text{-a.s.}}}\quad{\begin{tikzpicture}}$
where we have moved the $Y_{t}$ input to right to obtain a more convenient way of drawing the diagram. An application of Proposition A.24(b) lets us rewrite the diagram further to
[…]
since the inner conditional is precisely the definition of $\mathrm{BF}_{t-1}$. Note that this reasoning is still valid at $t=0$ if we take all objects and morphisms at negative times to be given by $I$ and $\mathrm{id}_{I}$, respectively. ∎
We will consider instances of the Bayes filter in the usual Markov categories of interest in the subsequent subsections and end here with a few general remarks.
Remark 3.3.
[…]
Remark 3.4.
What happens when we instantiate the Bayes filter in a parametric Markov category $\mathcal{C}_{W}$, in the style of Remarks 2.7 and 2.14? This is possible, as our constructions again apply to that case, and we thereby automatically obtain a Bayes filter with parameters. Can it be described more concretely?
Modulo swapping of inputs, the conditionals in a parametric Markov category $\mathcal{C}_{W}$ can be constructed as the conditionals in $\mathcal{C}$ of their representatives [14, Lemma 2.10]. Therefore, the Bayes filter associated with a hidden Markov model in a parametric Markov category is represented simply by the corresponding conditional in $\mathcal{C}$ [^13] Additionally, a.s.-equalities in $\mathcal{C}_{W}$ correspond to a.s.-equalities in $\mathcal{C}$, in such a way that the uniqueness criteria for the two conditionals coincide. This gives a very general formalization of the intuitive idea that the Bayes filter for a parametrized hidden Markov model should be the Bayes filter for each of its parameter values.
Remark 3.5.
A treatment of hidden Markov models and Bayesian filtering in the setting of Markov categories can also be found in a recent paper by Virgo [45] based on stochastic Mealy machines. This amounts to considering time-homogeneous hidden Markov models but generalized to the form $X\to Y\otimes X$, where the transition kernel and observation kernel is unified into one morphism as in (2.10), and in addition also allowing inputs [^14] These can be used to extend the framework to Markov decision processes, for instance., which we do not consider in this paper.
Virgo’s work adopts a perspective slightly different from ours. Fixing an observation space $Y$, he forms a category $\mathsf{Generator}\mathopen{}\mathclose{{}\left(Y}\right)$ having generalized hidden Markov models (in the above sense) with observation space $Y$ as objects and morphisms those maps between the state spaces which preserve the structure [45, Section 2.1]. Assuming that the Markov category in question is strongly representable, Virgo then extends Bayesian filtering to a functor out of the category $\mathsf{Generator}\mathopen{}\mathclose{{}\left(Y}\right)$. This can be seen as functorially converting a hidden Markov model into a machine that updates Bayesian priors. It is further shown that this Bayesian filtering functor is a right adjoint [45, Theorem 2.6]. While Virgo’s hidden Markov models are partly more restrictive (time-homogeneous) and partly more general than ours, his work is complementary to ours insofar as it is concerned with different aspects of the Bayes filter. For instance, Virgo does not prove any version of Proposition 3.2, focusing instead on the adjunction of [45, Theorem 2.6] and its consequences (such as constructing conjugate priors).
3.2 The instantiated Bayes filter
Following Proposition 3.2, the Bayes filter $\mathrm{BF}_{t}$ is given by a conditional involving the morphisms $\mathrm{BF}_{t-1}$, $f_{t}$ and $g_{t}$. However, this conditional is still difficult to work with in practice since it involves the whole input space $Y_{\mathopen{}\mathclose{{}\left[t}\right]}=Y_{0}\otimes\cdots\otimes Y_{t-1}$ containing all observations made. For this reason, we introduce the instantiated Bayes filter here, which specializes the Bayes filter to a fixed sequence of observations. This simplifies the recursive computation of the conditional (see Proposition 3.7) and recovers well-known filters like the Kalman filter in the Gaussian case (see Section 3.3.4). The price to pay is that one does not compute the whole filter $\mathrm{BF}_{t}$ but only its value on a particular sequence of inputs.
Here is the general categorical description of how this works.
Definition 3.6.
Given a sequence of deterministic morphisms
$y_{0}\colon I\to Y_{0},\qquad y_{1}\colon I\to Y_{1},\qquad\ldots,\qquad y_{t}\colon I\to Y_{t}$
up to time $t\in\mathopen{}\mathclose{{}\left\{0,\dots,n}\right\}$, which we abbreviate by $\mathbf{y}_{\mathopen{}\mathclose{{}\left[t}\right]}$, the instantiated Bayes filter is
[…]
The non-uniqueness of $\mathrm{BF}_{t}$ discussed before now has more severe consequences, as the instantiated Bayes filter is not well-defined in general. We will discuss this further in the case of $\mathsf{BorelStoch}$ in Section 3.3.2.
We have the following recursive formula for $\mathrm{IBF}_{t}$ analogous to Proposition 3.2.
Proposition 3.7.
The instantiated Bayes filter can be recursively computed through conditionals as
[…]
for all $t=0,\dots,n$, where we take $\mathrm{IBF}_{-1}\coloneqq\mathrm{id}_{I}$ to be the trivial morphism.
So explicitly, the $t=0$ case is given by
[…]
Each step of the recursion involves a choice of conditional, and making such choices throughout the recursion merely results in one version of the instantiated Bayes filter. We will discuss this further in the examples below.
Proof.
Proposition 3.2 is the recursion
[…]
This implies that
[…]
where we have used A.26 together with the definition of $\mathrm{IBF}_{t-1}(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t-1}\right]})$. ∎
The formula (3.5) allows us to compute our posterior distribution for the hidden state $x_{t}$ through the following two steps:
  1. (a)
    Prediction step: We use the posterior from our previous guess for the hidden state $x_{t-1}$ and apply the transition kernel $f_{t}$ in order to predict the next hidden state $x_{t}$.
  2. (b)
    Update step: We use the observation $y_{t}$ and the deterministic state $g_{t}$ to update our guess for the hidden state $x_{t}$ by conditioning on the observation at time $t$ being $y_{t}$.
This is the known procedure for the classical Bayes filter [39, Theorem 4.1], but now generalized to the categorical setting.
3.3 Examples of the instantiated Bayes filter
In the following, we show that the instantiated Bayes filter recovers known filters for suitable choices of Markov categories. This includes the discrete Bayes filter in the case of $\mathsf{FinStoch}$ as well as the Kalman filter in the case of $\mathsf{Gauss}$.
3.3.1 The instantiated Bayes filter in $\mathsf{FinStoch}$
We now consider the instantiated Bayes filter in the Markov category $\mathsf{FinStoch}$ explicitly, a procedure sometimes known as the forward algorithm151515see Remark 4.7.. As in 2.13, the hidden Markov model up to time $t\in\mathopen{}\mathclose{{}\left[n}\right]$ is a probability distribution on the product of the finite sets $\mathopen{}\mathclose{{}\left(X_{i}}\right)_{i=0}^{t}$ and $\mathopen{}\mathclose{{}\left(Y_{i}}\right)_{i=0}^{t}$. It is ‘‘categorical’’ in the sense of these sets being finite. The sequence of deterministic morphisms $\mathbf{y}_{\mathopen{}\mathclose{{}\left[t}\right]}$ corresponds to a sequence of elements $y_{0}\in Y_{0},\ldots,y_{t}\in Y_{t}$, to be thought of as samples from the distribution over observations $p_{\mathopen{}\mathclose{{}\left[t}\right]}^{Y}$. To then compute the instantiated Bayes filter $\mathrm{IBF}_{t}$ recursively, we consider the prediction and update steps as explained above separately. The prediction step produces the distribution $q_{t}\colon I\to X_{t}$ given by
$q_{t}\mathopen{}\mathclose{{}\left(x_{t}}\right)=\sum_{x_{t-1}}f_{t}\mathopen{}\mathclose{{}\left(x_{t}\,|\,x_{t-1}}\right)\cdot\mathrm{IBF}_{t-1}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t-1}\right]}}\right)\mathopen{}\mathclose{{}\left(x_{t-1}}\right)$ (3.7)
for $t>0$, and $q_{0}=f_{0}$ to get the recursion going. Performing the update step on top of that results in
$\mathrm{IBF}_{t}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t}\right]}}\right)\mathopen{}\mathclose{{}\left(x_{t}}\right)=\frac{g_{t}\mathopen{}\mathclose{{}\left(y_{t}\,|\,x_{t}}\right)\,q_{t}\mathopen{}\mathclose{{}\left(x_{t}}\right)}{\sum_{x^{\prime}_{t}}g_{t}\mathopen{}\mathclose{{}\left(y_{t}\,|\,x^{\prime}_{t}}\right)\,q_{t}\mathopen{}\mathclose{{}\left(x^{\prime}_{t}}\right)}.$ (3.8)
This equation agrees with the Bayes filter in its standard form for discrete variables [37, Page 9, Equation 1.25]. It is worth noting that the right-hand side is well-defined whenever the denominator is nonzero, while $\mathrm{IBF}_{t}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t}\right]}}\right)$ is an arbitrary distribution whenever the denominator is zero. This is not an issue in the present discrete case since the denominator is nonzero for all $y_{t}$ that can occur as part of a sequence of observations with nonzero probability.
3.3.2 The instantiated Bayes filter in $\mathsf{BorelStoch}$
This construction immediately generalizes to $\mathsf{BorelStoch}$. So now our $X_{i}$ and $Y_{i}$ are standard Borel spaces, and the $f_{i}\colon X_{i-1}\to X_{i}$ and $g_{i}\colon X_{i}\to Y_{i}$ are measurable Markov kernels. Then again, a sequence of observations for the instantiated Bayes filter is given by a sequence of elements $y_{0}\in Y_{0},\dots,y_{t-1}\in Y_{t-1}$. The prediction step constructs a probability measure $q_{t}\colon I\to X_{t}$ given by
$q_{t}\mathopen{}\mathclose{{}\left(\mathrm{d}x_{t}}\right)=\int_{x_{t-1}\in X_{t-1}}f_{t}\mathopen{}\mathclose{{}\left(\mathrm{d}x_{t}\,|\,x_{t-1}}\right)\cdot\mathrm{IBF}_{t-1}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t-1}\right]}}\right)\mathopen{}\mathclose{{}\left(\mathrm{d}x_{t-1}}\right),$
for $t>0$, which is the generalization of (3.7) in $\mathsf{BorelStoch}$, and with initial condition $q_{0}=f_{0}$. The update step is more subtle: every individual observation value $y_{t}$, or entire observation sequence $\mathbf{y}_{\mathopen{}\mathclose{{}\left[t}\right]}$, typically has probability zero, and it is a well-known fact that conditioning on events of probability zero is ill-defined. [^16] See e.g. the Borel–Kolmogorov paradox for a particularly striking example of this [20, Section 15.7], which clearly illustrates that this lack of uniqueness is not an issue with our formalism but rather an unavoidable complication in measure-theoretic probability quite generally. This, in particular, means that the instantiated Bayes filter $\mathrm{IBF}_{t}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t}\right]}}\right)$ is completely arbitrary for every sequence $\mathbf{y}_{\mathopen{}\mathclose{{}\left[t}\right]}$ with probability zero, which is the typical case when one has continuous variables. What can be considered more well-defined—as per the uniqueness of conditionals in $\mathsf{BorelStoch}$ up to almost sure equality—is the totality of values $\mathrm{IBF}_{t}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t}\right]}}\right)$ for all $\mathbf{y}_{\mathopen{}\mathclose{{}\left[t}\right]}$, which is our Bayes filter $\mathrm{BF}_{t}$ again.
With this in mind, we can explicate the update step: it is given by the formation of product regular conditional probabilities, which amounts to the implicit equation
$\int_{y_{t}\in T}\int_{x_{t}\in X_{t}}\mathrm{IBF}_{t}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t}\right]}}\right)\mathopen{}\mathclose{{}\left(S}\right)\,g_{t}\mathopen{}\mathclose{{}\left(\mathrm{d}y_{t}\,|\,x_{t}}\right)\,q\mathopen{}\mathclose{{}\left(\mathrm{d}x_{t}}\right)=\int_{x_{t}\in S}g_{t}\mathopen{}\mathclose{{}\left(T\,|\,x_{t}}\right)\,q\mathopen{}\mathclose{{}\left(\mathrm{d}x_{t}}\right)$
for all measurable $S\subseteq X_{t}$ and $T\subseteq Y_{t}$. This follows as an instance of [9, Example 11.3], and can be interpreted as follows: both sides represent the probability that the hidden state $x_{t}$ is in $S$ and that the observation $y_{t}$ is in $T$, and $\mathrm{IBF}_{t}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t}\right]}}\right)$ is constructed precisely such that this probability can be calculated as on the left-hand side.
[5, Proposition 3.1.4] gives a similar filtering formula for hidden Markov models of continuous variables with the additional restriction that the processes $g_{t}\mathopen{}\mathclose{{}\left(\mathrm{d}y_{t}\,|\,x_{t}}\right)$ need to be absolutely continuous for all $x_{t}$ with respect to a single measure.
3.3.3 The instantiated Bayes filter in $\mathsf{FinSetMulti}$
Next, we compute the Bayes filter in a ‘‘possibilistic’’ case, that of $\mathsf{FinSetMulti}$. Recall from A.22(e) that a conditional $f_{|X}$ of an arbitrary morphism $f\colon A\to X\otimes Y$ is given simply by
$f_{|X}\mathopen{}\mathclose{{}\left(y\,|\,a,x}\right)=f\mathopen{}\mathclose{{}\left(x,y\,|\,a}\right).$
Given then a hidden Markov model as in 2.13(ii), the prediction step for the instantiated Bayes filter again takes the form
$q_{t}\mathopen{}\mathclose{{}\left(x_{t}}\right)=\sum_{x_{t-1}}f_{t}\mathopen{}\mathclose{{}\left(x_{t}\,|\,x_{t-1}}\right)\cdot\mathrm{IBF}_{t-1}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t-1}\right]}}\right)\mathopen{}\mathclose{{}\left(x_{t-1}}\right).$
This now means that $x_{t}$ is considered possible if and only if $x_{t-1}$ is possible according to our previous estimate of the hidden state $\mathrm{IBF}_{t-1}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t-1}\right]}}\right)$ and $x_{t}$ is possible according to the transition kernel $f_{t}$. The update step is given simply by
$\mathrm{IBF}_{t}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t}\right]}}\right)\mathopen{}\mathclose{{}\left(x_{t}}\right)=g_{t}\mathopen{}\mathclose{{}\left(y_{t}\,|\,x_{t}}\right)\cdot q_{t}\mathopen{}\mathclose{{}\left(x_{t}}\right).$
Since this multiplication amounts to a logical conjunction of possibilities, this now means that the state $x_{t}$ is considered possible if and only if it was considered possible before and is in addition compatible with the observed $y_{t}$ according to the observation map $g_{t}$.
As far as we know, this has not been studied before in the context of automata theory. We call this version of the Bayes filter the possibilistic filter. [^17] We omit mention of Bayes here, since this version of the Bayes filter no longer uses Bayesian updating in the usual sense.
3.3.4 The Kalman filter as the instantiated Bayes filter
The well-known Kalman filter is the special case of the Bayes filter for Gaussian distributions. It was initially formulated as an optimization problem [23, 24] and was considered from a Bayesian perspective only later [17]. This re-derivation utilized a version of Bayes’ rule for probability density functions: the equations for the prediction and update steps were obtained from the parametrization of the probability densities for multivariate Gaussian distributions.
Here, we show that in the Markov category $\mathsf{Gauss}$, our instantiated Bayes filter turns into a generalized version of the Kalman filter. So as in 2.13(iii), the ${\mathopen{}\mathclose{{}\left(x_{t}}\right)}_{t=0}^{n}$ and ${\mathopen{}\mathclose{{}\left(y_{t}}\right)}_{t=0}^{n}$ are Gaussian vectors which are related via
$x_{t}=A_{t}x_{t-1}+\mathcal{N}\mathopen{}\mathclose{{}\left(v_{t},Q_{t}}\right),\qquad y_{t}=H_{t}x_{t}+\mathcal{N}\mathopen{}\mathclose{{}\left(w_{t},R_{t}}\right),$
where our notation loosely follows that of [39, Section 4.3]. In particular, we also write $m_{t}$ and $P_{t}$ for the mean and covariance matrix of the Gaussian distribution of $\mathrm{IBF}_{t}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t}\right]}}\right)$, so that
$\mathrm{IBF}_{t}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t}\right]}}\right)=\mathcal{N}\mathopen{}\mathclose{{}\left(m_{t},P_{t}}\right).$
This equation determines $\mathrm{IBF}_{t}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t}\right]}}\right)$ completely, and we may think of $m_{t}$ as our best guess of the hidden state given the observations up to time $t$, while $P_{t}$ is the covariance matrix of our uncertainty in this guess. The problem solved by the Kalman filter is the calculation of $m_{t}$ and $P_{t}$ recursively from $m_{t-1}$ and $P_{t-1}$ together with the new observation $y_{t}$. Note that we suppress the dependence on $\mathbf{y}_{\mathopen{}\mathclose{{}\left[t}\right]}$ in $m_{t}$ and $P_{t}$ from the notation.
For the prediction step, we simply apply the transition kernel $f_{t}$ to $\mathrm{IBF}_{t-1}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t-1}\right]}}\right)$. By the composition formula in $\mathsf{Gauss}$ (A.2), this produces a new Gaussian with parameters
$\tilde{m}_{t}=A_{t}m_{t-1}+v_{t},\qquad\tilde{P}_{t}=A_{t}P_{t-1}A_{t}^{t}+Q_{t},$ (3.9)
where the tildes remind us that this models the prediction obtained from the filter at time $t-1$ before taking the observation $y_{t}$ into account. The joint distribution with the expected observation $y_{t}$, or more formally the right-hand side of (3.5) before the conditioning, can then be calculated to be given by the Gaussian
$\begin{pmatrix}x_{t}\\ y_{t}\end{pmatrix}=\mathcal{N}\mathopen{}\mathclose{{}\left(a_{t},C_{t}}\right)$
with parameters
$a_{t}=\begin{pmatrix}\tilde{m}_{t}\\ H_{t}\tilde{m}_{t}+w_{t}\end{pmatrix}\qquad\text{ and }\qquad C_{t}=\begin{pmatrix}\tilde{P}_{t}&\tilde{P}_{t}H_{t}^{t}\\ H_{t}\tilde{P}_{t}&S_{t}\end{pmatrix},$
where
$S_{t}\coloneqq H_{t}\tilde{P}_{t}H_{t}^{t}+R_{t}$ (3.10)
is the so-called innovation covariance. Now $\mathrm{IBF}_{t}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t}\right]}}\right)$ is given by the conditional of this Gaussian with respect to $y_{t}$. Using the known formulas for conditionals in $\mathsf{Gauss}$ from A.22(d) and introducing the optimal Kalman gain [^18] As in A.22(d), we write $S_{t}^{-}$ for the Moore-Penrose pseudoinverse of $S$. In practice $S_{t}$ will usually be invertible, and then simply $S_{t}^{-}=S_{t}^{-1}$.
$K_{t}=\tilde{P}_{t}H_{t}^{t}S_{t}^{-}$ (3.11)
as shorthand notation, we obtain that
$\mathrm{IBF}_{t}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t}\right]}}\right)=\mathcal{N}\mathopen{}\mathclose{{}\left(m_{t},P_{t}}\right)$
with
$m_{t}=\tilde{m}_{t}+K_{t}\mathopen{}\mathclose{{}\left(y_{t}-H_{t}\tilde{m}_{t}-w_{t}}\right),\qquad P_{t}=\mathopen{}\mathclose{{}\left(I-K_{t}H_{t}}\right)\tilde{P}_{t}.$
Together with (3.9), (3.10), and (3.11), these formulas form a mild generalization of the standard Kalman filter formulas [39, Theorem 4.2]: the latter are recovered as the special case $v_{t}=w_{t}=0$, which amounts to assuming linearity and centered Gaussian noise for both the transition kernels and the observation kernels.
There are many other variations on the Kalman filter, some of which may fit well into this framework with the appropriate choice of Markov category. For example, some Kalman filter variations use different representations of uncertainty to improve robustness to numerical error from matrix inversions in the Bayesian inversion step [3]. Others account for different structures in the state and measurement spaces, such as filters over Riemannian manifolds [28]. Many variations attempt to handle nonlinear transition kernels and observation maps by approximating their state propagations [2, 22]. These approximating computations may be more challenging to neatly formalize categorically, but this fact may be a topic of interest for future research. While the details in these filters are highly varied from each other, the general predict-update structure is common to all of them, which indicates that the categorical Bayes filter can serve as a unifying abstract formulation for this structure.
Remark 3.8.
Remark 3.4 on the Bayes filter in parametric Markov categories has an instantiated analogue. Deterministic states in a parametric Markov category $\mathcal{C}_{W}$ correspond to deterministic morphisms in $\mathcal{C}$ with domain $W$.
Thus, the representative in $\mathcal{C}$ of an instantiated Bayes filter associated to a hidden Markov model and sequence of observations in $\mathcal{C}_{W}$ is given by precomposing the deterministic representatives of the observations and then feeding a copy of the parameter $W$ to each of these (as well as the parameter input of the representative of the Bayes filter).
As with the un-instantiated Bayes filter, the recursion of Proposition 3.7 can be translated into a statement in $\mathcal{C}$.
3.4 The Bayes filter recursion as a single string diagram
Our definition of the Bayes filter $\mathrm{BF}_{t}\colon Y_{\mathopen{}\mathclose{{}\left[t}\right]}\to X_{t}$ should be thought of as follows: given a sequence of observations as input, $\mathrm{BF}_{t}$ returns a sample from the posterior distribution of the hidden state. Although that distribution itself is precisely what is specified by $\mathrm{BF}_{t}$, it is not directly accessible as an output if one thinks of $\mathrm{BF}_{t}$ as a process that produces a random output. This situation changes in a representable Markov category (A.8), where $\mathrm{BF}_{t}$ has a deterministic counterpart $\mathrm{BF}_{t}^{\sharp}\colon Y_{\mathopen{}\mathclose{{}\left[t}\right]}\to PX_{t}$, which can be thought of as the actual function that returns the posterior distribution given a sequence of observations and which recovers $\mathrm{BF}_{t}$ after composition with the sampling morphism $\mathsf{samp}\colon PX_{t}\to X_{t}$.
Remark 3.9.
[…]
The goal of this subsection is to rewrite the recursion for the Bayes filter in terms of $\mathrm{BF}_{t}^{\sharp}$, which will let us encode the recursive construction into a single string diagram. Throughout, we assume that our Markov category $\mathcal{C}$ is a.s.-compatibly representable in addition to having conditionals, and we consider a hidden Markov model in the form of Proposition 2.8(iv) as before.
Notation 3.10.
For $t=0,\dots,n$, let $\mathrm{u}_{t}\colon PX_{t-1}\otimes Y_{t}\to X_{t}$ be the morphism
[…]
As before, this uses the convention that $X_{-1}=I$, and thus also $PX_{-1}=I$, so that $\mathrm{u}_{0}=\mathrm{BF}_{0}$. Here is how the morphisms $u_{t}$ act themselves as update maps for the Bayes filter.
Proposition 3.11.
We have
{tikzpicture}${\begin{tikzpicture}}\qquad=_{{p_{\mathopen{}\mathclose{{}\left[t}\right]}^{Y}\text{-a.s.}}}\qquad{\begin{tikzpicture}}$ (3.12)
for all $t=0,\dots,n$.
Proof.
This is a consequence of the equations where we have used Proposition 3.2 in the first step and A.26 together with the fact that $\mathrm{BF}_{n-1}^{\sharp}$ is deterministic in the third step. ∎
Corollary 3.12.
The deterministic counterpart of the Bayes filter satisfies the recursion
{tikzpicture}${\begin{tikzpicture}}\qquad=_{{p_{\mathopen{}\mathclose{{}\left[t}\right]}^{Y}\text{-a.s.}}}\qquad{\begin{tikzpicture}}$ (3.13)
Proof.
The right-hand side is deterministic as $\mathrm{u}_{t}^{\sharp}$ and $\mathrm{BF}_{t-1}^{\sharp}$ are, and deterministic morphisms form a symmetric monoidal subcategory. Post-composition with $\mathsf{samp}$ shows that it is the deterministic counterpart of the right-hand side of (3.12). Therefore the claim follows from (3.12) because we are in a a.s.-compatibly representable Markov category (A.11). ∎
Since the recursion of 3.12 is easy to expand, we immediately arrive at the following.
Corollary 3.13.
In any a.s.-compatibly representable Markov category with conditionals, the Bayes filter can be written as the single string diagram
{tikzpicture}${\begin{tikzpicture}}\qquad=_{{p_{\mathopen{}\mathclose{{}\left[t}\right]}^{Y}\text{-a.s.}}}\qquad{\begin{tikzpicture}}$
which furthermore recovers $\mathrm{BF}_{t}$ itself upon composition with $\mathsf{samp}$.
Remark 3.14.
If $\mathcal{C}$ is a.s.-compatibly representable, so is the parametric Markov category $\mathcal{C}_{W}$ for any object $W$ [14, Lemma 3.25]. Further, the deterministic counterpart of a morphism in $\mathcal{C}_{W}$ is represented in $\mathcal{C}$ by the deterministic counterpart of its representative in $\mathcal{C}$.
Consequently, as with Remark 3.4 and Remark 3.8, the results of this section also apply to $\mathcal{C}_{W}$. This immediately shows that 3.13 holds even in case all morphisms involved have an additional dependence on some parameter object $W$. This amounts to every box in our diagrams, including the $\mathrm{u}_{i}$, connecting to a copy of the $W$ input as in (2.6).
3.5 The Bayes filter as a Markov chain
Consider a hidden Markov model and the sequence of posterior distributions that describe an observer’s guess about the hidden state. Since the observations are generically random, the posterior distributions are also random. It is natural to consider them as a stochastic process in their own right and to ask whether they form a Markov chain. In this subsection, we show that this is indeed the case, generalizing [6, Lemma 2.4] to the categorical setting.
Throughout, we work in an a.s.-compatibly representable Markov category $\mathcal{C}$ with conditionals and consider a hidden Markov model and its Bayes filter as before. In these terms, the stochastic process of posterior distributions up to some time $t\in[n]$ is given schematically by the filter process
[…]
where the dashed wires indicate what the pattern is.
In order to spell this out more formally, let us introduce the following auxiliary morphisms.
Definition 3.15.
Let
$\lambda_{t}\colon Y_{0}\otimes\dots\otimes Y_{t}\longrightarrow PX_{0}\otimes\dots\otimes PX_{t}$
be defined recursively, starting with $\lambda_{0}\coloneqq\mathrm{BF}_{0}^{\sharp}$, and for $t=1,\dots,n$ by
[…]
As before, we could also consider the start of the recursion to be $\lambda_{-1}\coloneqq\mathrm{id}_{I}$ with $Y_{-1}=I$ and $PX_{-1}=X_{-1}=I$. In terms of these $\lambda_{t}$’s, the filter process is now formally defined as
[…]
Lemma 3.16.
For all $t\in\mathopen{}\mathclose{{}\left[n}\right]$, we have
[…]
Proof.
For $t=0$, this is simply the statement that $\lambda_{0}=\mathrm{BF}_{0}^{\sharp}$ is deterministic. The induction step follows similarly straightforwardly by the fact that $\mathrm{BF}_{t}^{\sharp}$ is deterministic. [^19] For a more abstract proof, note that all morphisms involved are deterministic, and it is therefore enough to show that all single-output marginals of the equation hold. We leave the details to the reader.
The following other sequence of auxiliary morphisms will describe the transition kernels of the filter process.
Definition 3.17.
For $t\in\mathopen{}\mathclose{{}\left[n}\right]$, let $h_{t}$ be the morphism
[…]
For $t=0$, we follow the convention $X_{-1}\coloneqq I$, so that $h_{0}=\mathrm{u}_{0}^{\sharp}g_{0}f_{0}$.
Remark 3.18.
Intuitively, $h_{t}$ takes a distribution on $X_{t-1}$, samples from it and determines a sample of the observation in $Y_{t}$, and then applies a Bayesian updating to the original distribution with respect to this observation. In $\mathsf{FinStoch}$, our $h_{t}^{\sharp}$ seems to match up with the ‘‘abstract hidden Markov models’’ of McIver et al. described at [27, Section 3.B] and [1, Section 14.3].
Before proceeding, it will be helpful to note that the process of observations can be generated recursively using the Bayes filter in the following sense.
Lemma 3.19.
For all $t\in\mathopen{}\mathclose{{}\left[n}\right]$, we have
[…]
Proof.
At $t=0$, this amounts to $p^{Y}_{\mathopen{}\mathclose{{}\left[0}\right]}=g_{0}f_{0}$, which is obvious. And with $t+1$ in place of $t$, the claim follows upon composing the definition of the Bayes filter (3.1) with $g_{t+1}f_{t+1}$. ∎
The following result now shows that the filter process is indeed a Markov chain with initial distribution $h_{0}$ and transition kernels $h_{t}$.
Theorem 3.20.
For all $t\in\mathopen{}\mathclose{{}\left[n}\right]$,
[…]
Proof.
We proceed by induction on $t$. The base case $t=0$ asserts $\mathrm{BF}_{0}^{\sharp}p^{Y}_{0}=h_{0}$, which holds because of $\mathrm{u}_{0}=\mathrm{BF}_{0}$ and $p^{Y}_{0}=g_{0}f_{0}$.
For the inductive step, we consider $t\geq 1$ such that the desired equation holds for $\mathrm{BFP}_{t-1}$, and we then show that it holds for $\mathrm{BFP}_{t}$. Expanding the definitions gives
[…]
Applying 3.19, this evaluates further to
[…]
where the second equation uses the recursion formula of 3.12. By the fact that $\mathrm{BF}_{t-1}^{\sharp}$ is the deterministic counterpart of $\mathrm{BF}_{t}$, we can also write this as
[…]
where the second equation holds by 3.16 and the definition of $h_{t}$. The claim now follows straightforwardly by the induction hypothesis. ∎
Remark 3.21.
This result may have some significance in the context of information security. Indeed, if the observations in the hidden Markov model are the observations of an adversary, then it is of interest to make statements about the adversary’s guess of the hidden state and how it evolves in time. The fact that the filter process is a Markov chain indicates that given the adversary’s guess at some time $t$, learning about their guess at an earlier time does not help in predicting their guess at a later time.
4 Smoothing in Markov categories
While the Bayes filter infers the most recent state on $X_{t}$ with respect to a history of measurements up to time $t$, it is sometimes desirable to go back and make inferences about the hidden state at an earlier time. So in this section, we consider observations up to time $n$ and the problem of inferring the hidden state at any time $t\in\mathopen{}\mathclose{{}\left[n}\right]$.
4.1 The Bayes smoother
To begin, analogously to 3.1 we define a smoother of a hidden Markov model as the conditional of some earlier state on a sequence of observations.
Definition 4.1.
For $p\colon I\to\mathopen{}\mathclose{{}\left(Y\otimes X}\right)_{\mathopen{}\mathclose{{}\left[n}\right]}$ a hidden Markov model, the Bayes smoother at time $t$ is the conditional
[…]
Working directly with this definition has similar practical limitations as with the Bayes filter: the size of the space that one conditions on grows exponentially with the number of observations $n$. Therefore, we again develop recursive formulas that make it easier to work with.
4.1.1 The forward-backward algorithm
As a start, note that constructing the entire joint distribution $p$ only to marginalize over all but one of the hidden state spaces seems wasteful. Using the structure of the hidden Markov model, we will see that this can indeed be avoided by factoring the relevant marginal into two separate parts and computing each part separately, roughly like so:
[…]
Then, using string diagram rewrites, we can reduce these morphisms to simpler forms and construct them recursively. The recursion to compute the factor on the left-hand side of the dashed line is known as the forward algorithm, and that for the right-hand side is known as the backward algorithm. These morphisms are then composed and conditioned on $Y_{0},\dots,Y_{n}$, a procedure called forward-backward algorithm.
These procedures are traditionally formulated concretely in $\mathsf{FinStoch}$ [33], and we now present the categorical generalization. As usual, this specializes also to a measure-theoretic formulation in $\mathsf{BorelStoch}$ and to a possibilistic formulation in $\mathsf{FinSetMulti}$. As we will see, the categorical formulation is so simple that it makes the algorithm seem almost trivial.
The goal of the forward algorithm is to compute the left part of (4.2), which is traditionally denoted by $\alpha$ [33, Section III].
Notation 4.2.
For a hidden Markov model $p$ as in 2.10 and $t\in\mathopen{}\mathclose{{}\left[n}\right]$, we write $\alpha_{t}\colon I\to Y_{\mathopen{}\mathclose{{}\left[t}\right]}\otimes X_{t}$ for the marginal given by
[…]
Remark 4.3.
By definition, the Bayes filter $\mathrm{BF}_{t}$ is the conditional of $\alpha_{t}$ on $Y_{\mathopen{}\mathclose{{}\left[t}\right]}$.
The naïve approach to computing $\alpha_{t}$ in practice would first calculate the entire $p_{\mathopen{}\mathclose{{}\left[t}\right]}\colon I\to\mathopen{}\mathclose{{}\left(Y\otimes X}\right)_{\mathopen{}\mathclose{{}\left[t}\right]}$ and then marginalize out $X_{\mathopen{}\mathclose{{}\left[t-1}\right]}$. But already in $\mathsf{FinStoch}$, this would involve many redundant calculations, resulting in a computation that quickly gets intractable in practice. Instead, the forward algorithm computes $\alpha_{t}$ by the following recursion that is linear in $t$.
Proposition 4.4 (The forward algorithm).
The morphisms $\alpha_{t}$ satisfy the recursion formula
[…]
for all $t\in\mathopen{}\mathclose{{}\left[n}\right]$, starting with $\alpha_{-1}\coloneqq\mathrm{id}_{I}$.
Proof.
Using the co-unitality of the comonoid structures, the right-hand side of (4.3) reduces to the form
[…]
The claimed recursion then follows from the obvious self-similarity of this string diagram. ∎
In contrast to the forward algorithm, the goal of the backward algorithm is to compute the right part of (4.2), traditionally denoted $\beta$ [33, Section III].
Notation 4.5.
For a hidden Markov model $p$ and $t\in\mathopen{}\mathclose{{}\left[n}\right]$, we write $\beta_{t}\colon X_{t}\to Y_{(t,n]}$ for
[…]
The naïve approach to computing this would be to construct the whole morphism $X_{t}\to\mathopen{}\mathclose{{}\left(Y\otimes X}\right)_{(t:n]}$ from the transition kernels and observation kernels and then marginalize out $X_{\mathopen{}\mathclose{{}\left[t-1}\right]}$, but this would again face exponential complexity. The following backward recursion formula constitutes the categorical version of the backward algorithm, which reduces this complexity to linear in $n-t$.
Proposition 4.6 (The backward algorithm).
The morphisms $\beta_{t}$ satisfy the backward recursion formula
[…]
for all $t\in\mathopen{}\mathclose{{}\left[n}\right]$, starting with $\beta_{n+1}\coloneqq\mathrm{id}_{I}$ where $X_{n+1}\coloneqq Y_{n+1}\coloneqq I$.
More explicitly, the first step of the recursion constructs $\beta_{n}$ as
[…]
which can alternatively be taken to define the base case of the recursion.
Proof.
This simple proof is analogous to the proof of Proposition 4.4. We rewrite the right-hand side of (4.6) using co-unitality of copy,
[…]
where we keep the right-most branch with the deletion on $X_{n}$ since that corresponds to $\beta_{n}$. The self-similarity of this string diagram then yields the claimed recursion. ∎
The Bayes smoother $\mathrm{BS}^{n}_{t}$ from 4.1 can now be written as
[…]
since by (4.2), the expression inside the dashed box is precisely the relevant marginal of the whole hidden Markov model $p$. The forward-backward algorithm now consists of using the recursions of Propositions 4.4 and 4.6 to compute the morphism inside the dashed box.
Remark 4.7.
There are some differences in the literature concerning the definitions of the forward and the forward-backward algorithm, and in particular on whether some conditioning on observations is already considered part of the algorithm or not. Common to almost all prior treatments is that they take place only in $\mathsf{FinStoch}$ only, although Murphy [30, Section 18.3.2.3] does discuss an equivalent for continuous variables with density. Otherwise accounts differ in the details. For example, Rabiner [33] defines $\alpha_{t}$ as we do, and conditioning on observations is only done in subsequent steps, such as trying to estimate the most likely hidden state (known as the Baum–Welch algorithm). Murphy [30, Sections 17.4.2, 17.4.3] however defines $\alpha_{t}$ to coincide with our definition of the Bayes filter $\mathrm{BF}_{t}$, and their description of the forward algorithm is more akin to the instantiated Bayes filter in $\mathsf{FinStoch}$ as we have in Section 3.3.1. Subsequently, their definition of the forward-backward algorithm corresponds to a $\mathsf{FinStoch}$ implementation of our Equation 4.17 below.
For clarity, we use $\alpha_{t}$ and $\beta_{t}$ to denote quantities that are not conditioned, and we reserve notation such as $\mathrm{BF}_{t}$ and $\mathrm{BS}^{n}_{t}$ for morphisms that have been conditioned on observations and thus perform Bayesian inference.
The formation of the conditional in (4.9) can still pose challenges. Let us thus note that the conditioning on $Y_{\mathopen{}\mathclose{{}\left[t}\right]}$ in (4.9) can be rewritten in terms of the Bayes filter.
Lemma 4.8.
The Bayes smoother can also be expressed as
[…]
Proof.
Using Proposition A.24, we can rewrite (4.9) as
[…]
Now by definition of the Bayes filter, the inner dashed box is precisely $\mathrm{BF}_{t}$. ∎
The conditioning on $Y_{(t,n]}$ in (4.10) may still pose challenges if $n-t$ is not small. A solution to this has been explored in $\mathsf{Gauss}$ called the two-filter smoother [25, 39], but we instead explore an alternative smoothing algorithm below.
4.1.2 Fixed-interval smoothing
Here, we develop the categorical version of an existing recursion formula for the Bayes smoothers $\mathrm{BS}^{n}_{t}$ in terms of the Bayes filters $\mathrm{BF}_{t}$. This is known as fixed-interval smoothing [2, Section 8.6], in reference to the time interval for the observation sequence being fixed. As with the recursion formulas for the Bayes filter considered in Section 3, this reduces the complexity to linear in $n$ and $t$.
Proposition 4.9 (Fixed-interval smoother).
The following backwards recursion formula for $\mathrm{BS}^{n}_{t}$ holds true for all $t\in\mathopen{}\mathclose{{}\left[0,n}\right]$:
[…]
where $X_{n+1}=I$, meaning that $\mathrm{BS}^{n}_{n+1}=\operatorname{\mathrm{del}}_{Y_{\mathopen{}\mathclose{{}\left[n}\right]}}$ and $f_{n+1}=\operatorname{\mathrm{del}}_{X_{n}}$.
More explicitly, the initial $t=n$ step of the recursion constructs $\mathrm{BS}^{n}_{n}=\mathrm{BF}_{n}$, which can alternatively be taken to be the start of the recursion.
Proof.
Using 4.8 and the backwards procedure from Proposition 4.6, we can express $\mathrm{BS}^{n}_{t}$ as
[…]
We take the parametric Bayesian inverse of $f_{t+1}$ from A.29, expressing it as
[…]
where we brought $f_{t+1,\mathrm{BF}_{t}}^{\dagger}$ outside the box and nested conditionals using Proposition A.24 and A.26. The inner box equals $\mathrm{BF}_{t+1}$ by Proposition 3.2, resulting in
[…]
which is equal to the stated expression in the proposition by 4.8 and A.29.
Finally, the correctness of the initial condition at $t=n$ is trivial since $\mathrm{BS}^{n}_{n}=\mathrm{BF}_{n}$ by definition. ∎
4.2 The instantiated Bayes smoother
As with the Bayes filter, working with the uninstantiated smoother $\mathrm{BS}^{n}_{t}$ can be difficult in practice already because the input space grows exponentially with $n$. Here, we introduce the instantiated Bayes smoother by specializing the Bayes smoother to a fixed sequence of observations, much in the same fashion as we had done for the Bayes filter in Section 3.2, which is often enough in practice. From here we can easily construct an instantiated version of the fixed-interval smoothing recursion, which in the Gaussian case recovers the well-known Rauch-Tung-Striebel smoother.
Definition 4.10.
Given a sequence of deterministic morphisms
$y_{0}\colon I\to Y_{0},\qquad y_{1}\colon I\to Y_{1},\qquad\ldots,\qquad y_{n}\colon I\to Y_{n}$
which we abbreviate by $\mathbf{y}_{\mathopen{}\mathclose{{}\left[n}\right]}$, the instantiated Bayes smoother is
[…]
Combining this definition with (4.9) now produces the instantiated version of the forward-backward algorithm,
[…]
where the $\alpha_{t}$ and $\beta_{t}$ are computed recursively as per Propositions 4.4 and 4.6.
We can subsequently use A.26 to produce the instantiated version of 4.8:
[…]
Similarly as with the instantiated Bayes filter, we also have an instantiated version of the fixed-interval smoothing recursion, which is how smoothing is typically performed in practice in combination with the recursion for the instantiated Bayes filter from Proposition 3.7. This changes the size of the conditioned channel from $Y_{(t,n]}$ to $X_{t+1}$.
Proposition 4.11 (Instantiated fixed-interval smoother).
The instantiated Bayes smoother can be computed through the following backwards recursion:
[…]
for all $t\in[0,n)$, where the recursion starts at $\mathrm{IBS}^{n}_{n}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[n}\right]}}\right)=\mathrm{IBF}_{n}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[n}\right]}}\right)$.
We omit the proof since it is straightforward and very similar to Proposition 3.7, utilizing A.6 and A.26 on Proposition 4.9.
4.3 Examples of the instantiated Bayes smoother
We briefly consider what form the forward-backward algorithm and the fixed-interval smoother take in particular Markov categories, limiting our presentation to $\mathsf{FinStoch}$ for the former and $\mathsf{Gauss}$ for the latter.
4.3.1 The forward-backward algorithm in $\mathsf{FinStoch}$
Traditionally, the forward-backward algorithm is employed mostly in $\mathsf{FinStoch}$. The forward procedure is usually taken to start at $t=0$ and then takes the following form, which is a straightforward translation of (4.4) and (4.7) into $\mathsf{FinStoch}$.
The forward procedure is:
  1. (a)
    Initialize as $\alpha_{0}\mathopen{}\mathclose{{}\left(y_{0},x_{0}}\right)\coloneqq g_{0}\mathopen{}\mathclose{{}\left(y_{0}\,|\,x_{0}}\right)\,f_{0}\mathopen{}\mathclose{{}\left(x_{0}}\right)$.
  2. (b)
    For $t=1,\ldots,n$, compute
    $\alpha_{t}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t}\right]},x_{t}}\right)=g_{t}\mathopen{}\mathclose{{}\left(y_{t}\,|\,x_{t}}\right)\sum_{x_{t-1}}f_{t}\mathopen{}\mathclose{{}\left(x_{t}\,|\,x_{t-1}}\right)\,\alpha_{t-1}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t-1}\right]},x_{t-1}}\right)$ (4.19)
The backward procedure is:
  1. (a)
    Initialize as $\beta_{n}\mathopen{}\mathclose{{}\left(x_{n}}\right)=1\quad\forall x_{n}\in X_{n}$. [^20] This $\beta_{n}\colon X_{n}\to I$ is sometimes described as a “uniform distribution” which is technically correct as the only distribution on $I$ is uniform, but we stress that this is not a uniform distribution on $X_{n}$, already because $X_{n}$ is the domain rather than the codomain of $\beta_{n}$.
  2. (b)
    For $t=n-1,\dots,0$, compute
    $\beta_{t}\mathopen{}\mathclose{{}\left(\mathbf{y}_{(t,n]}\,|\,x_{t}}\right)=\sum_{x_{t+1}}g_{t+1}\mathopen{}\mathclose{{}\left(y_{t+1}\,|\,x_{t+1}}\right)\,\beta_{t+1}\mathopen{}\mathclose{{}\left(\mathbf{y}_{(t+1,n]}\,|\,x_{t+1}}\right)\,f_{t+1}\mathopen{}\mathclose{{}\left(x_{t+1}|x_{t}}\right).$
Finally, performing the conditioning with the observed values $\mathbf{y}_{[n]}$ plugged in as in (4.16) produces the desired result,
$\mathrm{IBS}^{n}_{t}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[n}\right]}}\right)\mathopen{}\mathclose{{}\left(x_{t}}\right)=\frac{\alpha_{t}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t}\right]},x_{t}}\right)\beta_{t}\mathopen{}\mathclose{{}\left(\mathbf{y}_{(t,n]}\,|\,x_{t}}\right)}{\sum_{x_{t}^{\prime}}\alpha_{t}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t}\right]},x_{t}^{\prime}}\right)\beta_{t}\mathopen{}\mathclose{{}\left(\mathbf{y}_{(t,n]}\,|\,x_{t}^{\prime}}\right)}$ (4.20)
We leave the instantiation of this algorithm in other Markov categories, like $\mathsf{BorelStoch}$, $\mathsf{Gauss}$, $\mathsf{FinSetMulti}$ and parametric Markov categories, to the reader.
4.3.2 The RTS smoother as the instantiated Bayesian smoother
Shortly after the Kalman filter, a Gaussian fixed-interval backwards recursion was developed [34], now commonly known as the Rauch-Tung-Striebel (RTS) smoother or Kalman smoother [30, Section 18.3.2]. Similar to the Kalman filter in Section 3.3.4, we show that our categorical fixed-interval smoother recursion recovers a slightly generalized version of the RTS smoother with biased noise when calculated in $\mathsf{Gauss}$.
Keeping notation as in the Kalman filter case from Section 3.3.4, we additionally write the parameters for the filter output as
$\mathrm{IBF}_{t}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t}\right]}}\right)=\mathcal{N}\mathopen{}\mathclose{{}\left(m_{t},P_{t}}\right),$
noting that we assume these quantities to have been computed already e.g. through the Kalman filter. To indicate the relation, let us denote the output parameters of the instantiated smoother with ‘‘hats,’’ meaning that
$\mathrm{IBS}^{n}_{t}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[n}\right]}}\right)=\mathcal{N}\mathopen{}\mathclose{{}\left(\hat{m}_{t},\hat{P}_{t}}\right).$
To determine the recursion formulas for these new parameters, we first write out the joint state inside the dashed box in (4.18) as
$\begin{pmatrix}x_{t}\\ \tilde{x}_{t+1}\end{pmatrix}=\mathcal{N}(a_{t},S_{t})$
with parameters
$a_{t}=\begin{pmatrix}m_{t}\\ \tilde{m}_{t+1}\end{pmatrix}\qquad\text{ and }\qquad S_{t}=\begin{pmatrix}P_{t}&P_{t}A_{t+1}^{t}\\ A_{t+1}P_{t}&\tilde{P}_{t+1}\end{pmatrix}$
determined by
$\tilde{m}_{t+1}=A_{t+1}m_{t}+v_{t+1},\qquad\tilde{P}_{t+1}=A_{t+1}P_{t}A_{t+1}^{t}+Q_{t+1},$ (4.21)
as in (3.9) for the Kalman filter. The conditional of this joint state is represented by the equation
$\hat{x}_{t}=C_{t}\hat{x}_{t+1}+\mathcal{N}\mathopen{}\mathclose{{}\left(\ m_{t}-C_{t}\tilde{m}_{t+1},\ P_{t}-C_{t}\tilde{P}_{t+1}C_{t}^{t}}\right)$
where we have introduced the optimal smoother gain
$C_{t}\coloneqq P_{t}A_{t+1}^{t}\tilde{P}_{t+1}^{-}.$
Composing this with $\mathrm{IBS}^{n}_{t+1}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[n}\right]}}\right)$, we obtain the recursion formulas that define the RTS smoother,
$\hat{m}_{t}=m_{t}+C_{t}\mathopen{}\mathclose{{}\left(\hat{m}_{t+1}-\tilde{m}_{t+1}}\right),\qquad\hat{P}_{t}=P_{t}+C_{t}\mathopen{}\mathclose{{}\left(\hat{P}_{t+1}-\tilde{P}_{t+1}}\right)C_{t}^{t}.$ (4.22)
where the tilde quantities are defined by (4.21).
Appendix A Background on Markov categories and conditional independence
In this section, we provide some basic background on Markov categories, marginalization, infinite tensor products, conditionals and conditional independence. Relative to our prior works, the only new content here is the dashed box notation for conditionals introduced in A.23, for which A.26 is a new result.
Notation A.1.
Throughout, we consider a symmetric monoidal category $\mathcal{C}$ with monoidal unit $I$. We will use standard string diagram notation for morphisms in $\mathcal{C}$ with the convention that the source of a morphism is at the bottom and the target on top.
Given a distinguished comonoid structure given by a comultiplication $X\to X\otimes X$ and counit $X\to I$ on an object $X\in\mathcal{C}$, we use the following symbols in our string diagrams to denote these two maps:
[…]
A.1 Markov categories
The following definition of Markov category was introduced under this name in [9, Definition 2.1], but had already appeared earlier as affine CD category in [7, Definition 2.3]. [^21] We refer to [11, Remark 2.2] for a more detailed account of the history.
Definition A.2.
A semicartesian [^22] Recall that this means that the monoidal unit is terminal. symmetric monoidal category $\mathcal{C}$ is a Markov category if every object $X$ carries a distinguished commutative comonoid structure [^23] The relevant piece of data here is $\operatorname{\mathrm{copy}}_{X}$, as $\operatorname{\mathrm{del}}_{X}$ must be the unique morphism to the terminal object.
$\operatorname{\mathrm{copy}}_{X}\colon X\to X\otimes X,\qquad\operatorname{\mathrm{del}}_{X}\colon X\to I,$
which is compatible with the tensor product in the sense that for all objects $X$ and $Y$, the diagram
[…]
commutes.
In terms of A.1, the commutativity condition (A.1) can be written more intuitively as
[…]
Example A.3.
The ‘‘trivial’’ examples of Markov categories are the cartesian monoidal categories (cf. [9, Remark 2.4]). The comonoid structures here are simply given by the diagonal maps as copy morphisms.
Example A.4.
In the following, we present several pertinent examples of Markov categories. For a detailed treatment of these examples, we refer to [9].
  1. (a)
    $\mathsf{FinStoch}$ is the Markov category with finite sets as objects and stochastic matrices ${\mathopen{}\mathclose{{}\left(f\mathopen{}\mathclose{{}\left(y\,|\,x}\right)}\right)}_{x\in X,y\in Y}$ as morphisms, with matrix multiplication as composition. The symmetric monoidal structure is given by the Cartesian product of sets and the tensor product (Kronecker product) of matrices. This Markov category models discrete probability theory with finite sets. In particular, a morphism $p\colon I\to X$ is a probability distribution on a finite set $X$.
  2. (b)
    $\mathsf{Stoch}$ is the Markov category with arbitrary measurable spaces as objects and measurable Markov kernels as morphisms. The symmetric monoidal structure is given by the usual product of measurable spaces and the tensor product of Markov kernels. This Markov category models general measure-theoretic probability. In particular, a morphism $p\colon I\to X$ is just a probability measure on a measurable space $X$.
  3. (c)
    $\mathsf{BorelStoch}$ is given by $\mathsf{Stoch}$ restricted to standard Borel spaces as objects. This is the Markov category that we usually use for measure-theoretic probability, since it is better-behaved than $\mathsf{Stoch}$ and since standard Borel spaces are sufficient for most applications.
  4. (d)
    $\mathsf{Gauss}$ is the Markov category modelling Gaussian distributions and kernels. Its objects are the Euclidean spaces $\mathbb{R}^{n}$, morphisms $p\colon I\to\mathbb{R}^{n}$ are Gaussian probability measures, and general morphisms $f\colon\mathbb{R}^{n}\to\mathbb{R}^{m}$ can be understood as stochastic maps of the form
    $x\longmapsto Ax+\xi,$
    where $A$ is any real $m\times n$ matrix and $\xi$ is an independent Gaussian variable specified by its mean $\mu$ and covariance matrix $\Sigma$. We will also write
    $x\mapsto Ax+\mathcal{N}\mathopen{}\mathclose{{}\left(\mu,\Sigma}\right)$
    in order to denote this map. Two such maps
    $x\mapsto Ax+\mathcal{N}\mathopen{}\mathclose{{}\left(\mu,\Sigma}\right),\qquad y\mapsto By+\mathcal{N}\mathopen{}\mathclose{{}\left(\nu,\Lambda}\right)$
    compose to
    $\begin{split}x\mapsto{}&{}\mathopen{}\mathclose{{}\left(BA}\right)x+B\mathcal{N}\mathopen{}\mathclose{{}\left(\mu,\Sigma}\right)+\mathcal{N}\mathopen{}\mathclose{{}\left(\nu,\Lambda}\right)\\ &=\mathopen{}\mathclose{{}\left(BA}\right)x+\mathcal{N}\mathopen{}\mathclose{{}\left(B\mu+\nu,B\Sigma B^{t}+\Lambda}\right),\end{split}$ (A.2)
    where the last equation holds as a consequence of the assumption that the two noise terms are independent.
  5. (e)
    $\mathsf{FinSetMulti}$ is the Markov category of finite sets and multivalued maps, which models nondeterminism in the computer science sense. A morphism $f\colon X\to Y$ is given by a matrix ${\mathopen{}\mathclose{{}\left(f\mathopen{}\mathclose{{}\left(y\,|\,x}\right)}\right)}_{y\in Y,x\in X}$ with entries in $\mathopen{}\mathclose{{}\left\{0,1}\right\}$, subject to the condition that for every $x$ we have $f\mathopen{}\mathclose{{}\left(y\,|\,x}\right)=1$ for some $y$. We interpret the case $f\mathopen{}\mathclose{{}\left(y\,|\,x}\right)=1$ as saying that output $y$ is possible on input $x$, while $f\mathopen{}\mathclose{{}\left(y\,|\,x}\right)=0$ says that $y$ is impossible on input $x$. We define the composition of two morphisms $f\colon X\to Y$ and $g\colon Y\to Z$ via
    $\mathopen{}\mathclose{{}\left(g\circ f}\right)\mathopen{}\mathclose{{}\left(z\,|\,x}\right)\coloneqq\sum_{y\in Y}g\mathopen{}\mathclose{{}\left(z\,|\,y}\right)\cdot f\mathopen{}\mathclose{{}\left(y\,|\,x}\right),$
    where we use the convention $1+1=1$. This is formally the same composition as in the case of $\mathsf{FinStoch}$, just with different arithmetic.
Almost sure equality is a core concept in probability theory, and as such it must have a counterpart in categorical probability theory.
Definition A.5 ([7, Definition 5.1][9, Definition 13.1]).
Consider morphisms $p\colon A\to X$ and $f,g\colon X\to Y$ in a Markov category. We say that $f$ and $g$ are $p$-almost surely equal, denoted $f=_{{p\text{-a.s.}}}g$, if
[…]
Intuitively this is thought of as $f$ and $g$ behaving the same on the support of $p$, where $p$ is itself thought of as a state depending on a parameter $A$.
We may think of a general morphism in a Markov category as a process that takes some input and produces a typically random output. However, the following subclass of morphisms are those that we should think of as genuine functions whose output is not random.
Definition A.6 ([9, Definition 10.1]).
A morphism $f$ in a Markov category $\mathcal{C}$ is deterministic if it commutes with copying, that is if
[…]
The wide [^24] The adjective wide means that it contains all objects of $\mathcal{C}$. subcategory comprised of the deterministic morphisms is called the deterministic subcategory $\mathcal{C}_{\det}$. This can be shown to be a cartesian monoidal subcategory of $\mathcal{C}$ [9, Remark 10.13]. For example $\mathsf{FinStoch}_{\det}$ is the category of finite sets and functions, $\mathsf{BorelStoch}_{\det}$ is the category of standard Borel spaces and measurable functions, $\mathsf{Gauss}_{\det}$ is the category of Euclidean spaces and affine maps, and $\mathsf{FinSetMulti}_{\det}$ is again the category of finite sets and functions.
Notation A.7.
For a morphism of type
[…]
the marginal of $f$ on $X$ is given by
[…]
We may speak complementarily of marginalizing out $Y$.
Similarly, we can talk about the marginal of $f$ on $Y$ or of marginalizing out $X$. We can also marginalize out several outputs, which is equivalent to marginalizing out their tensor product.
A.2 Representability
One $\mathsf{BorelStoch}$, one can associate to an object $X$ another object $PX$, namely the space of distributions on $X$ equipped with a suitable $\sigma$-algebra, such that measurable Markov kernels $A\to X$ correspond to measurable functions $A\to PX$. Representability axiomatizes this phenomenon in the general setting of Markov categories, providing a sense in which objects have associated ‘‘distribution objects’’ within the same Markov category.
Definition A.8 ([14, Definition 3.10]).
A Markov category $\mathcal{C}$ is representable if the inclusion functor $\mathcal{C}_{\det}\rightarrowtail\mathcal{C}$ has a right adjoint $P\colon\mathcal{C}\to\mathcal{C}_{\det}$.
This definition can be rephrased in terms of the counit of the adjunction, which we denote
$\mathsf{samp}\colon P\to\mathrm{id}_{\mathcal{C}}.$
Indeed by general abstract nonsense, the existence of the adjoint $P$ is equivalent to the existence of, for each object $X$, a distribution object $PX$ and a sampling map
$\mathsf{samp}_{X}\colon PX\to X$
such that every morphism $f\colon A\to PX$ factors into a unique deterministic morphism $f^{\sharp}\colon A\to PX$ followed by $\mathsf{samp}_{X}$. In terms of string diagrams, this is
[…]
Notation A.9.
In a representable Markov category, we write $f^{\sharp}$ for the deterministic counterpart of a morphism $f$. We will also often drop the component index from $\mathsf{samp}$ for brevity, as it can be inferred from context.
Examples A.10.
We briefly review the representability of two of our recurring examples. The details of many others can be found in [14].
  1. (a)
    $\mathsf{BorelStoch}$ is representable, with right adjoint $P$ sending a space $X$ to the space of distributions on it. The idea is that the deterministic counterpart of a morphism $f\colon X\to Y$ is the measurable map $f\colon X\to PY$ that sends every $x\in X$ to the distribution $f\mathopen{}\mathclose{{}\left(-\,|\,x}\right)$.
  2. (b)
    $\mathsf{FinSetMulti}$ is representable, with right adjoint $P$ sending a set $X$ to $2^{X}\setminus\mathopen{}\mathclose{{}\left\{\emptyset}\right\}$. The idea is that the deterministic counterpart of a morphism $f\colon X\to Y$ is the function $f\colon X\to PY$ that sends every $x\in X$ to the subset $\mathopen{}\mathclose{{}\left\{y\in Y\,|\,f\mathopen{}\mathclose{{}\left(y\,|\,x}\right)=1}\right\}\subseteq Y$ of possible outcomes.
A slightly stronger notion of representability is often useful.
Definition A.11.
A representable Markov category is called a.s.-compatibly representable if for all $p\colon A\to X$ and $f,g\colon X\to Y$,
$f=_{{p\text{-a.s.}}}g\quad\implies\quad f^{\sharp}=_{{p\text{-a.s.}}}g^{\sharp}$ (A.4)
Remark A.12.
The converse of this implication holds for representable Markov categories in general, as can be seen by post-composing with the relevant sampling map.
Examples A.13.
Most representable Markov categories we know of are also a.s.-compatibly representable. In particular, this is the case for $\mathsf{BorelStoch}$ [14, Example 3.19]. Also, $\mathsf{FinSetMulti}$ is a.s.-compatibly representable. A direct proof of this claim is straightforward (see [13, Proposition A.4.4 and Example A.4.6]).
An example of a representable Markov category that is not a.s.-compatibly representable is given in [14, Section 3.3].
Much more can be said about representable Markov categories, for instance regarding the interaction of $P$ with the monoidal structure. We do not need these explicitly in this paper, but refer the interested reader to [14] for more details.
A.3 Kolmogorov products
In the main text, considering morphisms whose codomain is a product $X_{1}\otimes\dots\otimes X_{t}$ is one ingredient for the joint distribution of a Markov chain up to time $t$. In order to talk about the joint distribution over all infinitely many times, we need a notion of infinite tensor product.
Notation A.14.
For a finite family of objects $\mathopen{}\mathclose{{}\left(X_{i}}\right)_{i\in F}$ of objects in a symmetric monoidal category, we write
$X_{F}\coloneqq\bigotimes_{i\in F}X_{i},$
where the symmetry allows us to leave the order of the factors unspecified.
Definition A.15 ([12, Definition 4.1]).
For a family ${\mathopen{}\mathclose{{}\left(X_{i}}\right)}_{i\in I}$ of objects in a Markov category $\mathcal{C}$, a Kolmogorov product is an object $X_{I}$ together with a family of deterministic morphisms
$\mathopen{}\mathclose{{}\left(\pi_{F}\colon X_{I}\to X_{F}}\right)_{F\subseteq I\text{ finite}}$
such that for all objects $A$ and $Y$, composition with the $\pi_{F}\otimes\mathrm{id}_{Y}$ establishes a bijection
$\mathcal{C}\mathopen{}\mathclose{{}\left(A,X_{I}\otimes Y}\right)\cong\mathopen{}\mathclose{{}\left\{\mathrm{compatible}\text{ }\mathrm{families}\>\mathopen{}\mathclose{{}\left(g_{F}\colon A\to X_{F}\otimes Y}\right)_{F\subseteq I\>\mathrm{finite}}}\right\},$
where a family $\mathopen{}\mathclose{{}\left(g_{F}\colon A\to X_{F}\otimes Y}\right)_{F\subseteq I\text{ finite}}$ is compatible if for all $F\subseteq F^{\prime}$ finite, $g_{F^{\prime}}$ marginalizes to $g_{F}$.
Taking $Y=I$ shows that $X_{I}$ is in particular a limit of the diagram formed by the finite powers $X_{F}$ with the marginalizations as arrows in the diagram. It follows that $X_{I}$ is unique up to unique isomorphism, and we thus may also speak of the Kolmogorov product of a family of objects.
Remark A.16.
The primary inspiration for the definition of Kolmogorov product is Kolmogorov’s extension theorem, which states that probability measures on a countable product correspond (say for standard Borel spaces) to marginalization-compatible families of probability measures on the finite subproducts. This entails that countably infinite Kolmogorov products exist in $\mathsf{BorelStoch}$ and are given by the usual product of measurable spaces.
Proposition A.17 ([12, Proposition 4.3]).
Kolmogorov products in a Markov category $\mathcal{C}$ are categorical products in $\mathcal{C}_{\det}$.
A.4 Conditionals and conditioning
We now turn to another central concept that we use throughout the paper.
Definition A.18 ([9, Definition 11.5]).
Given a morphism $f\colon A\to X\otimes Y$ in a Markov category $\mathcal{C}$, a morphism $f_{|Y}\colon A\otimes Y\to X$ is a conditional of $f$ with respect to $Y$ if the following equation holds:
[…]
We say that $\mathcal{C}$ has conditionals if every such a conditionals always exists.
Remark A.19.
For a morphism $p\colon I\to X\otimes Y$ with trivial input, the defining equation of a conditional $p_{|Y}$ simplifies to
[…]
This is an equation going back to [7, Section 3], and specializes to the standard defining equation of a conditional distribution in $\mathsf{FinStoch}$,
$p\mathopen{}\mathclose{{}\left(x,y}\right)=p_{|Y}\mathopen{}\mathclose{{}\left(x\,|\,y}\right)\,p\mathopen{}\mathclose{{}\left(y}\right).$
Remark A.20.
Conditionals are not, strictly speaking, unique. [^25] This is indeed already the case in $\mathsf{FinStoch}$. Even for $p\colon I\to X\otimes Y$ in $\mathsf{FinStoch}$, the usual formula $p_{|Y}\mathopen{}\mathclose{{}\left(x\,|\,y}\right)=\frac{p\mathopen{}\mathclose{{}\left(x,y}\right)}{p\mathopen{}\mathclose{{}\left(y}\right)}$ is ill-defined as soon as $p\mathopen{}\mathclose{{}\left(y}\right)=0$. According to A.18, the correct defining equation in general is $p\mathopen{}\mathclose{{}\left(x,y}\right)=p_{|Y}\mathopen{}\mathclose{{}\left(x\,|\,y}\right)\cdot p\mathopen{}\mathclose{{}\left(y}\right)$, and we see that $p_{|Y}\mathopen{}\mathclose{{}\left(x\,|\,y}\right)$ is completely arbitrary for all $y$ with $p\mathopen{}\mathclose{{}\left(y}\right)=0$. However, by pre-composing the defining equation (A.5) with $\operatorname{\mathrm{copy}}_{A}$, it can be shown that the conditional $f_{Y}$ is unique up to almost sure equality with respect to
[…]
Due to this ‘‘almost uniqueness’’, we may also speak of ‘‘the’’ conditional, as long as it is ensured that all statements under consideration are valid regardless of the particular choice of conditional.
Remark A.21.
In an a.s.-compatibly representable Markov category, also the deterministic counterpart $f_{|Y}^{\sharp}$ is unique up to almost sure equality with respect to the same morphism (A.6). This follows from the almost uniqueness of $f_{|Y}$ by directly applying the definition (A.4).
Example A.22.
In the following, we discuss the existence of conditionals in the Markov categories given in A.4.
  1. (a)
    $\mathsf{FinStoch}$ has conditionals. The conditional $f_{|Y}$ of a morphism $f\colon A\to X\otimes Y$ is given by
    $f_{|Y}\mathopen{}\mathclose{{}\left(x\,|\,y,a}\right)\coloneqq\frac{f\mathopen{}\mathclose{{}\left(x,y\,|\,a}\right)}{f\mathopen{}\mathclose{{}\left(y\,|\,a}\right)}$
    where the denominator is the marginal
    $f\mathopen{}\mathclose{{}\left(y\,|\,a}\right)=\sum_{x\in X}f\mathopen{}\mathclose{{}\left(x,y\,|\,a}\right),$
    and we address the problem of $\frac{0}{0}$ by taking $f_{|Y}\mathopen{}\mathclose{{}\left(x\,|\,y,a}\right)$ to be an arbitrary distribution on $X$ whenever $f\mathopen{}\mathclose{{}\left(y\,|\,a}\right)=0$.
  2. (b)
    In $\mathsf{Stoch}$, a conditional $p_{|Y}\colon Y\to X$ for a probability measure on a product space $p\colon I\to X\otimes Y$ is usually called regular conditional distribution. More generally, for a given Markov kernel $f\colon A\to X\otimes Y$, a conditional is a Markov kernel $f_{|Y}\colon A\otimes Y\to X$ such that
    $f\mathopen{}\mathclose{{}\left(S\times T\,|\,a}\right)=\int_{y\in T}f_{|Y}\mathopen{}\mathclose{{}\left(S\,|\,y,a}\right)\cdot f\mathopen{}\mathclose{{}\left(\mathrm{d}y\,|\,a}\right)$
    for all measurable $S\subseteq X$ and $T\subseteq Y$ and all $a\in A$. Indeed since two probability measures on a product measurable space are equal if and only if they are equal on products of measurable sets, this equation is a complete translation of (A.5) into measure-theoretic language. Such a conditional does not exist in general [9, Example 11.3].
  3. (c)
    $\mathsf{BorelStoch}$ has conditionals, since every standard Borel space has regular conditional distributions, and these can be formed even for distributions with measurable dependence on an additional parameter $a\in A$ [9, Example 11.3].
  4. (d)
    $\mathsf{Gauss}$ has conditionals. A morphism $f\colon A\to X\otimes Y$ is given by
    $x=Ma+\xi\quad\text{ and }\quad y=Na+\eta,$
    where $M$ and $N$ are suitably sized matrices and $\xi$ and $\eta$ are (not necessarily independent) Gaussian random variables. Equivalently, we can write
    $\begin{pmatrix}x\\ y\end{pmatrix}=\begin{pmatrix}M\\ N\end{pmatrix}a+\mathcal{N}\mathopen{}\mathclose{{}\left(v,C}\right)$
    where the Gaussian noise has mean $v$ and covariance matrix $C$ of the form
    $v=\begin{pmatrix}s\\ t\end{pmatrix}\quad\text{ and }\quad C=\begin{pmatrix}C_{\xi\xi}&C_{\xi\eta}\\ C_{\eta\xi}&C_{\eta\eta}\end{pmatrix}.$
    As per [9, Example 11.8], a conditional $f_{|Y}\colon A\otimes Y\to X$ is given by
    $x=\mathopen{}\mathclose{{}\left(N-C_{\xi\eta}C_{\eta\eta}^{-}M}\right)a+C_{\xi\eta}C_{\eta\eta}^{-}y+\tau$
    where $C_{\eta\eta}^{-}$ is the Moore–Penrose pseudoinverse of $C_{\eta\eta}$, and the Gaussian noise $\tau$ is specified by
    $\mathbf{E}\mathopen{}\mathclose{{}\left[\tau}\right]=s-C_{\xi\eta}C_{\eta\eta}^{-}t\quad\text{ and }\quad\mathbf{Var}\mathopen{}\mathclose{{}\left[\tau}\right]=C_{\xi\xi}-C_{\xi\eta}C_{\eta\eta}^{-}C_{\eta\xi}.$
    See also [43, Section 6.5] for a more geometric construction of conditionals in a closely related category of extended Gaussians.
  5. (e)
    $\mathsf{FinSetMulti}$ has conditionals [10, Proposition 16]. For a given morphism $f\colon A\to X\otimes Y$, a conditional $f_{|Y}$ is simply given by
    $f_{|Y}\mathopen{}\mathclose{{}\left(x\,|\,y,a}\right)=f\mathopen{}\mathclose{{}\left(x,y\,|\,a}\right).$
Notation A.23.
Following Jacobs [18, Section 7.5/6], we denote a conditional as in A.18 also by
[…]
In this notation, the defining equation of the conditional reads as
[…]
Although we will use this dashed box notation for conditionals throughout the paper, it is important to keep in mind the non-uniqueness of conditionals as discussed in Remark A.20. As noted there, the conditional is only unique up to almost sure equality, and this uniqueness is often sufficient for the composite morphism depicted by a string diagram to be well-defined. However, a dashed box as a standalone part of a string diagram is usually not well-defined.
The following coherence results for conditionals, originally proven as [9, Lemmas 11.11-12], admit particularly elegant depictions in terms of the dashed box notation [18, Exercise 7.5.4].
Proposition A.24.
In any Markov category $\mathcal{C}$:
  1. (a)
    If $f\colon A\to X\otimes Y\otimes Z$ admits a conditional $f_{|Z}$, and if $f_{|Z}$ has a further conditional $(f_{|Z})_{|Y}$, then this ‘‘double conditional’’ is also a conditional $f_{|\mathopen{}\mathclose{{}\left(Z\otimes Y}\right)}$,
[…]
    where the almost sure equality is with respect to
[…]
  2. (b)
    For $f\colon B\to W\otimes Y$ having a conditional $f_{|Y}$ and any $g\colon A\otimes W\to X$, we have
[…]
    where the almost sure equality is with respect to
[…]
Remark A.25.
The second item shows in particular that conditioning commutes with marginalization: for any $f\colon A\to X\otimes Y\otimes Z$ for which the conditional $f_{|Z}$ exists, we have
[…]
More symbolically, this equation takes the form
$f_{X|Z}=\mathopen{}\mathclose{{}\left(f_{|Z}}\right)_{X}.$
In particular, it is a consequence that the conditional $f_{X|Z}$ exists.
We have a similar and new coherence result for precomposition, but unlike the previous coherence results it involves a determinism assumption.
Lemma A.26.
For any $f\colon A\otimes B\to X\otimes Y$ such that $f_{|Y}$ exists and any deterministic $g\colon C\to B$, we have
[…]
where the almost sure equality is with respect to
[…]
Proof.
We prove the statement for $A=I$, since then the general case follows by using $\mathrm{id}_{A}\otimes g$ in place of $g$. In this case we get
[…]
where we have used the definition of the conditional of $f$ in the first step and the assumption that $g$ is deterministic in the second step. But this shows the claim that $f_{|Y}\circ\mathopen{}\mathclose{{}\left(g\otimes\mathrm{id}_{Y}}\right)$ is a conditional of $f\circ g$ with respect to $Y$. ∎
Remark A.27.
Assuming $g$ to be deterministic in A.26 is necessary, in the sense that if the conclusion holds for $A=I$ and $f=\operatorname{\mathrm{copy}}_{B}$, then $g$ is deterministic.
Indeed $\operatorname{\mathrm{copy}}_{B}$ has a conditional with respect to the second output given by $\mathrm{id}_{B}\otimes\operatorname{\mathrm{del}}_{B}$. Thus, if the equality of A.26 holds for $f\coloneqq\operatorname{\mathrm{copy}}_{B}$, it is the assertion that $g\otimes\operatorname{\mathrm{del}}_{B}$ is a conditional,
[…]
The defining equation for that conditional is
[…]
which states precisely that $g$ is deterministic.
Definition A.28 ([7, Definition 3.5]).
Consider a morphism $f\colon X\to Y$ and $p\colon I\to X$. Then a Bayesian inverse $f_{p}^{\dagger}$ of $f$ with respect to $p$ is a conditional
[…]
Expanding definitions, this is equivalent to $f^{\dagger}_{p}$ satisfying the equation
[…]
As with general conditionals, $f^{\dagger}_{p}$ is not strictly unique, but is rather defined up to almost sure equality with respect to $f\circ p$.
We can generalize this definition to the case where $p$ is not just a state but a general morphism. We use the same notation in terms of the dagger. [^26] Note that this dagger is not a dagger functor in the same sense as the one of vanilla Bayesian inversion is [9, Remark 13.10].
Definition A.29.
For morphisms $f\colon X\to Y$ and $g\colon A\to X$, a parametric Bayesian inverse $f_{g}^{\dagger}$ of $f$ with respect to $g$ is a conditional
[…]
Expanding definitions, this is equivalent to $f^{\dagger}_{g}$ satisfying the equation
[…]
Note that there is an equivalent mirror-image definition, where conditioning is done on the right and the above equation goes the other way. These two definitions are interchangeable by a swap, and whichever one is used should be obvious in context throughout the paper.
A.5 Conditional independence
In Markov categories with conditionals, there is a particularly well-behaved theory of conditional independence. We recall part of this here for the case of morphisms with trivial input, and refer to [10] for a general treatment.
Proposition A.30 ([7, Proposition 6.9]).
In a Markov category $\mathcal{C}$ with conditionals, for any $p\colon I\to X\otimes Y\otimes Z$ the following are equivalent:
  1. (a)
    We have [^27] In $\mathsf{FinStoch}$ for instance, this is the equation $p\mathopen{}\mathclose{{}\left(x,y,z}\right)=p_{|Y}\mathopen{}\mathclose{{}\left(x\,|\,y}\right)\,p_{|Y}\mathopen{}\mathclose{{}\left(z\,|\,y}\right)\,p\mathopen{}\mathclose{{}\left(y}\right).$
[…]
  2. (b)
    There is a morphism $f\colon Y\to X$ such that [^28] In $\mathsf{FinStoch}$, this is $p\mathopen{}\mathclose{{}\left(x,y,z}\right)=f\mathopen{}\mathclose{{}\left(x\,|\,y}\right)\,p\mathopen{}\mathclose{{}\left(y,z}\right),$ so that necessarily $f=p_{|Y}$.
[…]
If these hold, then the morphism $f$ is necessarily a conditional,
[…]
We reproduce the simple proof in our notation for convenience.
Proof.
Property (b) immediately implies the final claim upon marginalizing out $Z$. Using this and the existence of conditionals, property (b) can also be written as
[…]
as was to be shown. The converse implication from (a) to (b) holds by the same argument used in the other direction, with $f$ defined as the relevant conditional. ∎
Definition A.31.
If the equivalent properties of Proposition A.30 hold, we say that $p$ displays the conditional independence
$X\perp Z\>\mid\>Y.$
For a state with additional outputs, like $p\colon I\to W\otimes X\otimes Y\otimes Z$, we say that $p$ displays this conditional independence if marginalizing out $W$ results in a morphism $p_{XYZ}$ which satisfies the above definition.
Remark A.32.
Using properties such as the symmetry in $X$ and $Y$, one can produce further equivalent forms of conditional independence.